### v1

Abstract
Private data is typically larger and of higher quality than public data, offering great potential to improve LLM. However, its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments. To address this, the transformer-based split learning model has emerged, offloading most model parameters to the server while retaining only the embedding and output layers on clients to ensure privacy. However, it still faces significant challenges in security, efficiency, and adaptability: 1) embedding gradients are vulnerable to attacks, leading to reverse engineering of private data; 2) the autoregressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) fixed partition points lack adaptability to downstream tasks. In this paper, we introduce FL-LLaMA, a secure, efficient, and adaptive federated split framework based on LLaMA2. First, we place some input and output blocks on the local client and inject Gaussian noise into forward-pass hidden states, enabling secure end-to-end propagation. Second, we employ client-batch and server-hierarchical strategies to achieve parallel training, along with attention-mask compression and KV cache mechanisms to accelerate inference, reducing communication costs effectively. Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements and hardware limitations. Experiments on NLU, summarization and conversational QA tasks show that FL-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 2x train speedups and 8× inference speedups. Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FL-LLaMA in security and adaptability.
keywords:
Federated Learning , Large Language Model , Privacy Protection , Communication Efficiency , Split Learning
PACS:
07.05.Mh , 84.35.+i
MSC:
[2010] 68Q85 , 68W15 , 68T05 , 68T50 , 68W10
journal: Artificial Intelligence
[…]
{graphicalabstract}[Uncaptioned image]
[…]
{highlights}
We introduce the federated splitting LLM, FL-LLaMA. FL-LLaMA holds some input and output blocks on the local client and injects Gaussian noise into forward-pass hidden states, enabling secure end-to-end propagation.
We employ client-batch and server-hierarchical strategies to achieve parallel training, along with attention-mask compression and KV cache mechanisms to accelerate training and inference.
FL-LLaMA allows users to dynamically adjust the partition points for input/output blocks based on specific task requirements and hardware limitations.
1 Introduction
The success of Large Language Models (LLMs) can be attributed to their ability to leverage large amounts of public data and intricate parameter networks [1]. By training on extensive corpora, LLMs can capture statistical patterns in language, semantic structures, and subtle contextual relationships. However, private data holds a significant advantage over public datasets in terms of both quality and utibility. This is because private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies [2]. Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios [3]. Therefore, effectively incorporating private data will become one of the critical factors in further advancing the performance of LLMs.
Figure 1: Model architecture of FedBert and FL-LLaMA.
The distributed nature of private data across various silos and the significant computational demands of LLMs present substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices [4], enterprises servers [5] or medical institutions [6, 7], as centralized storage and processing of such data may raise privacy concerns and regulatory risks [8]. Federated Learning (FL) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation, but traditional FL frameworks, such as FedAvg [9] and FedProx [10], are infeasible for LLMs training scenarios. LLMs typically have billions of parameters, and their training process requires extensive computational resources, placing substantial demands on device memory and network bandwidth. This makes training the full model on ordinary client devices practically infeasible. Additionally, the communication overhead and heterogeneity inherent in LLMs training become even more pronounced within federated environments, making it challenging for models to efficiently converge or maintain generalization capabilities. As a result, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open and cutting-edge research topic.
As an alternative method, transformer-based split learning methods, named FedBERT [11], have been proposed to adapt LLMs, offloading most model parameters reside to the server while retaining only the embedding and output layers on client to ensure privacy. As shown in Fig. 1(a), FedBERT divides Transformers into three parts, i.e., embedding layers and output layers on client, while main Transformer layers on server. The forward propagation(FP) path is client-to-server-to-client. After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client. The core idea of FedBERT is to divide the calculation into two parts: the client devices is responsible for initial feature extraction from raw data and calculate final loss, while the server handles deeper inference and learning. Although FedBERT provides a potentially viable solution for resource-constrained clients scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the embedding layer are transmitted, which can be exploited to reconstruct original input data through beam search [12] or reverse engineering [13]. 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
In this paper, we introduce FL-LLaMA, a secure, efficient, and adaptive federated split framework based on LLaMA2, as shown in Fig. 1(b). To Enable Secure Propagation, we partition the transformers of LLaMA2 into three parts: input and output blocks on clients, while the remaining large parameters on server. During training, clients perform forward propagation to generate hidden states, add Gaussian noise for them, and send these noised states to the server for further processing. Through both gradients analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages. To Reduce Communication Cost, we employ client-batch and server-hierarchical strategies to achieve parallel training: the former batches data from different clients, while the latter leverages concurrent sub-server to enhance training efficiency. Durning inference, we introduce the attention-mask compressio and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of autoregressive inference. By synchronously maintaining cache states on both client and server, the originally cross-device transmition is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B). To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources. This adaptability enables optimal balance between performance and efficiency.
Experiments on NLU, summarization, and conversational QA tasks show that FL-LLaMA maintains performance comparable to centralized LLaMA2, and achieving up to 2x training and 8× inference speedups. Through inference acceleration, the average token generation time of FL-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the length of context.  [^2] Our code is available at: https://github.com/TAP-LLM/SplitFedLLM.
The innovations in this paper are as follows:
  • 1.
    We design a federated split framework tailored for resource-constrained clients scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
  • 2.
    We propose client-batch and server-hierarchical optimizations to enhance split learning efficiency, and use collaborative KV cache with attention mask compression to reduce autoregressive inference communication.
  • 3.
    We enables dynamic partition points based on downstream tasks and device capabilities, optimizing the trade-off between computational efficiency and model performance.
  • 4.
    Experimental results on NLU, summarization and conversation QA tasks show that our FL-LLaMA achieves performance comparable to centralized LLaMA2, and achieve up to 2x train speedups and 8× inference speedups.
This paper is an extension of the work presented in  [14]. There are five major differences from the original version: (i) the base model and fine-tuning method have been updated, from p-tuning v2 on ChatGLM in the original version to LoRA on LLaMA2, demonstrating the generalizability of our federated split framework; (ii) Gaussian noise is introduced during the forward pass of split learning to further enhance the security of the federated split LLMs; (iii) the use of attention-mask compression and collaborative KV cache mechanism significantly improves inference efficiency in the federated split setting; (iv) dynamic partitioning is supported, allowing users to customize split points via configuration files, thus improving adaptability; (v) the experimental evaluation has been extended by assessing FL-LLaMA on new summarization and conversational QA tasks, showcasing its versatility.
The rest of this paper is organized as follows. Section 2 reviews related work, and Section 3 presents the details of the federated split framework. Section 4 reports the experimental evaluation. Finally, Section 5 concludes the paper.
2 Related work
2.1 Fedrated Learning for LLMs
Federated Learning (FL) has emerged as a promising paradigm for training language model in a decentralized manner while preserving user privacy and ensuring data security. FedAvg [15] is one of the earliest and most widely adopted optimization algorithms, which has been successfully applied in multiple language modeling tasks [16, 17, 9]. In this approach, each client locally trains a model on its private data and sends the updates to a central server, which performs a weighted aggregation to produce a new global model. To reduce the number of local training rounds and accelerate convergence, Stremmel et al. [9] propose initializing the global model with a pretrained language model before applying FedAvg, which has proven effective in speeding up training. Ji et al. [18] introduce FedAtt, which incorporates a hierarchical soft attention mechanism during aggregation to adaptively weighs each client’s model parameters based on their contribution, improving both generalization and local adaptation.
In terms of LLMs, traditional FL frameworks face growing challenges due to increased communication and computational costs. To address the issue of massive model sizes, some federated LLMs framework  [19, 20, 3] has been proposed to leverage instructions tuning, Adapters and LoRA for local fine-tuning, which require updating only a small subset of parameters and significantly reducing communication costs while maintaining performance. But these methods still require the transmission of parameter quantities far exceeding the system’s carrying capacity. To further minimize communication overhead, gradient compression [21, 22] and quantization [23] mechanisms have been explored, which effectively reduce the volume of data exchanged, making them well-suited for federated fine-tuning of LLMs. Liu et al. [24] propose structural pruning approaches that update only selected parameter layers of LLMs based on Fisher Information. By narrowing the update scope, these methods effectively reduce the size of transmitted models without sacrificing overall performance.
Besides knowledge distillation methods between large and small language models [25, 26, 27] are proposed to avoid the high communication cost of transferring LLMs parameters directly. It utilize the smaller proxy models acting as intermediaries, facilitating knowledge transfer between clients and the server with substantially lower communication. Additionally, client heterogeneity remains a central challenge in FL, due to variations in computational capacity, data distributions, and task types across clients [28]. Bai et al. [29] and Wang et al. [30] introduce LoRA-based stacking strategies to balance the differences in client capabilities. They help ensure that models from clients with varying ranks can still be aggregated effectively, improving training stability and fairness.
While recent works of federated LLMs explores efficient fine-tuning and parameter reduction methods, resource-constrained clients still struggle with independent fine-tuning, especially with added compression overhead. Therefore, split learning has emerged as a viable alternative for such scenarios.
2.2 Split Learning for LLMs
Split learning [31] is gaining traction as a decentralized and collaborative training paradigm tailored for environments with limited computational resources [32, 33, 34, 35], particularly suited for scenarios involving mobile platforms or lightweight clients that lack access to GPUs. The core idea involves partitioning a deep neural network into separate sub-network, with each sub-network assigned to a different device for distributed processing.
To address the training of sequential data in LLMs, FedBERT [11] introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. In this approach, the BERT model is divided into three core components: the Embedding layer, the Transformer layer, and the Output layer. This separation allows the heavy Transformer layer to be trained on a centralized, high-performance server, while the lighter Embedding and Output layers are handled locally by client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points and susceptibility to privacy threats like embedding gradient attacks.
Recently, Su et al. [36] extend the scope of split learning by exploring scenarios with heterogeneous resource configurations, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time for LLMs. Furthermore, the repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities.
Our work diverges from these studies by targeting a scenario with a powerful server and resource-constrained clients holding private data. We prioritize enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
[…]
2.3 Attacks and Defenses
In federated learning, although it allows multiple clients to jointly fine-tune models without revealing raw data, there remains a risk of embedding-based reconstruction attacks, LoRA-based gradient inversion attacks and split learning-based data reconstructing attacks.
Attacks from embedding gradients have proven highly effective in reconstructing users’ private data. Gupta et al. [12] demonstrated that by identifying non-zero elements in the embedding gradients, an adversary can infer which tokens were used by the client, using beam search to reconstruct private data. To mitigate such attacks, they proposed freezing the embedding layer during training. Zhu et al. [37] briefly discussed defensive strategies such as injecting differential privacy noise [38, 39] or applying gradient clipping by zeroing out small gradients. Huang et al. [40] introduced a MixUp-based data augmentation method on the [CLS] token in BERT, while Yaldiz et al. [41] suggested applying cosine similarity checks on client-uploaded weights to detect and filter out malicious clients. But these defenses often degrade the model’s accuracy [42, 43].
In addition to embedding-based reconstruction attacks, Federated LLMs are also vulnerable to gradient inversion attacks via LoRA, due to frequent parameter sharing across clients [44, 45]. To address this, Liu et al. [46] propose DP-LoRA, which adds Gaussian noise to LoRA updates to ensure differential privacy [47] [39]. However, the combination of heterogeneous data, DP noise, and LoRA’s sensitivity to hyperparameters can lead to instability and performance loss. To overcome this, Sun et al. [48] introduced FFA-LoRA, which initializes the non-zero entries of the LoRA $A$ matrix from a Gaussian distribution and keeps matrix $A$ frozen while only fine-tuning the zero-initialized LoRA $B$ matrix, thus preventing the injected noise under differential privacy from being amplified. However, exposure of intermediate activations and gradients in distributed training still enables potential data reconstruction, prompting increased focus on secure split learning methods.
In split learning, Ege et al. [49] propose inferring labels by minimizing gradient differences between a pseudo-model and the actual model, achieving strong performance in classification tasks. However, this approach is impractical for LLMs due to the high computational cost of handling vocabularies exceeding 100,000 tokens, making real-time federated training infeasible. In image recognition, Dario et al. [50] show that a malicious server could reconstruct client data by exploiting hidden states via a Feature Space Hijacking Attack (FSHA). However, FSHA relies on access to a shadow dataset from the same domain and task—an unrealistic assumption in privacy-preserving settings, where servers lack access to such data.
Building on prior attacks, we explore a more severe threat model where a malicious client colludes with an honest-but-curious server by sharing its data, enabling the server to train an attack model that reconstructs plaintext from other clients using hidden states. Importantly, the server does not alter model outputs or gradients.
3 Framework
In this section, we present a detailed description of the FL-LLaMA framework, as shown in Fig. 2, which comprises five components: split learning, privacy preservation, parallel training, inference acceleration and adaptive partitioning. First, we adopt split learning to divide the LLaMA into three parts: the local client stores the first ${p}$ blocks and the last ${q}$ blocks, while the remaining ${k}$ blocks, comprising the majority of the model parameters, are hosted on the server. Next, we design a Gaussian noise injection mechanism applied to the forward-pass hidden states to ensure data privacy during transmission between the client and server. Third, we enhance the training efficiency of split learning through a client-batch and server-hierarchical parallelisms. Fourth, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms. Finally, the framework allows clients to adaptively adjust the split configuration based on their privacy preferences and available computational resources.
3.1 Split Learning
Suppossing that the LLMs comprises $N$ transformer blocks, the proposed FL-LLaMA framework partitions the LLMs into three segments: Part I contains the embedding layer and the first $p$ transformer blocks, denoted as $Blocks_{A}$; Part II includes the middle $k$ transformer blocks, denoted as $Blocks_{B}$; Part III consists of the last $q$ transformer blocks along with the output layer, denoted as $Blocks_{C}$, where $p+k+q=N$. For generative tasks, the model predicts the next token autoregressively, with data labels closely linked to the input text sequences. To safeguard client data privacy, the FL-LLaMA framework places Part I $Blocks_{A}$ and Part III $Blocks_{C}$ on the client side, while Part II $Blocks_{B}$ resides on the server side, thereby preventing direct exposure of client data. Such an architecture optimizes GPU memory allocation, as the server handles most of the model’s GPU memory burden, significantly reducing memory pressure on the client.
The architecture diagram of the proposed model is shown on the left of Fig. 2. Taking a single-client distributed training scenario as an example, at the beginning of training, the client utilizes local data to perform forward propagation through $Blocks_{A}$, resulting in hidden states $h_{A}$:
$h_{A}=f_{1\sim p}\circ f_{\text{embed}}(\mathbf{x}),$ (1)
where $f_{\text{embed}}(\cdot)$ denotes the embedding layer of the LLMs, and $f_{1\sim p}(\cdot)$ represents the sequential forward propagation through transformer block 1 to block $p$ at the client side.
After obtaining $h_{A}$, the client transmits it together with the position index and the attention mask matrix $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ to the server, where $b$ is the batch size and $seq\_len$ is the maximum context length in LLMs. Subsequently, $Blocks_{B}$ on the server side takes the received hidden states $h_{A}$ and attention mask $A_{mask}$ as inputs, performs computationally intensive tasks, and returns the resulting hidden states $h_{B}$ back to the client:
$h_{B}=f_{p+1\sim p+k}(h_{A},A_{mask}),$ (2)
where $f_{p+1\sim p+k}(\cdot)$ denotes sequential forward propagation through transformer blocks $p+1$ to $p+k$ on the server side.
Finally, $Blocks_{C}$ on the client-side performs computations to obtain the logits $\mathbf{z}$, computes the loss with local data, and initiates gradient backpropagation. These steps are defined by the following equations:
$\mathbf{z}=f_{p+k+1\sim N}(h_{B},A_{mask}),$ (3)
$\mathcal{L}=\mathcal{F}_{loss}(f_{out}(\mathbf{z}),\mathbf{y}),$ (4)
where $f_{p+k+1\sim N}(\cdot)$ denotes sequential forward propagation through transformer blocks $p+k+1$ to $N$ on the client side, $f_{out}(\cdot)$ is the output layer of the model, $\mathbf{y}$ is the targeting sentences and $\mathcal{F}_{loss}(\cdot)$ denotes the cross-entropy loss function.
After computing the loss, the client performs gradient backpropagation locally to update $Blocks_{C}$, then transmits the gradient of the hidden states $h_{B}$ back to the server. The server executes gradient backpropagation on $Blocks_{B}$, updates its parameters, and subsequently transmits the gradients of hidden states $h_{A}$ back to the client. Lastly, $Blocks_{A}$ performs gradient backpropagation, updates the model and embedding parameters, and completes a sequential training round.
3.2 Privacy Preservation
In distributed training of LLMs, the split learning paradigm partitions the entire model into multiple sub-models deployed across different physical nodes. As a result, the hidden states and gradients must be transferred across nodes via the network. Unlike traditional federated learning, which only shares gradients, this approach exposes mid-layer activations and their gradients, greatly increasing the risk of adversaries reverse-engineering user data and challenging the data protection framework of federated split LLMs.
[…]
3.2.1 Threat Model Defination
We design an extreme attack scenario, inspired by Dario et al. [50], to evaluate the robustness of FL-LLaMA. Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Fig. 3, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without their awareness.
Concretely, we consider an adversary that has full knowledge of the client-side model architecture and the number of layers deployed on the client, but does not have access to the actual model parameters. Thus, the honest-but-curious server locally initializes a decoder with the same architecture and depth as the client’s model, same with LLaMA2 configuration, and uses it as the attack model. Without loss of generality, we consider a two-client FL-LLaMA setting, where the server receives hidden states from both $client_{0}$ and $client_{1}$. During each training step, the server first performs a normal forward pass through $Blocks_{B}$ to obtain the intermediate hidden states $h_{B}$. It then feeds the hidden states from the malicious $client_{0}$ into the attack model and uses the corresponding private data $D_{\text{priv}}$ as supervision. Afterward, $h_{B}$ is split and sent back to the respective clients, allowing training to proceed as usual without disrupting the federated process.
Figure 4: Inject noise to hidden state $h_{1}$. Noise added to $h_{1}$ eventually protects the gradients of parameters.
3.2.2 Defense with Gaussion Noise Injection
We explore two defense strategies to protect the transmission in FL-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states [^3] It is noteworthy that we chose not to use the standard DPSGD [51], which adds noise after computing all gradients, because FL-LLaMA requires noise protection during both forward and backward passes, unlike DPSGD’s focus on post-backpropagation gradients.. The first strategy introduces random Gaussian noise into the initial gradient during backpropagation (i.e., the gradient $\nabla h_{B}$), while the second applies Gaussian noise to the client’s first forward-pass output (i.e., the hidden state $h_{A}$). Empirical results reveal that federated split LLMs are highly sensitive to gradient perturbations, so injecting noise into gradients often leads to instability in federated training, with the loss frequently diverging to NaN after several training steps. In contrast, perturbing the hidden state $h_{A}$ during the forward pass yields more stable training. Therefore, FL-LLaMA adopts the latter to enhance robustness while preserving privacy.
Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation. Fig. 4 illustrates the schematic of noise injection during the forward propagation of LLMs. Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer and $W_{n+1}$ for the weight of output linear layer. In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
$h_{2}$ $=\mathcal{A}(h_{1})+h_{1},$ (5)
$h_{3}$ $=\text{RMSNorm}(h_{2}),$ (6)
$h_{4}$ $=\mathcal{M}(h_{3})+h_{2},$ (7)
$=W_{n}[Silu(gate_{proj}(h_{3}))\cdot up_{proj}(h_{3})]+h_{2},$ (8)
$h_{5}$ $=W_{n+1}\cdot\text{RMSNorm}(h_{4}),$ (9)
$\hat{y}$ $=\text{softmax}(h_{5}),$ (10)
$\mathcal{L}$ $=CrossEntropy(y,\hat{y}),$ (11)
where $\text{RMSNorm}(*)$ is the layernorm function, $\mathcal{A}$ is the attention layer, $\mathcal{M}$ is the MLP layer.
During backpropagation, according to [52], the gradient of $W_{n}$ is:
$\frac{\partial L}{\partial W_{n}}$ $=\frac{\partial L}{\partial h_{5}}\cdot\frac{\partial h_{5}}{\partial h_{4}}\cdot\frac{\partial h_{4}}{\partial W_{n}},$ (12)
$\frac{\partial L}{\partial h_{5}}$ $=(\hat{y}-y),$ (13)
$\frac{\partial h_{5}}{\partial h_{4}}$ $=W_{n+1}\cdot J_{\text{RMSNorm}}(h_{4}),$ (14)
$\frac{\partial h_{4}}{\partial W_{n}}$ $=[Silu(gate_{proj}(h_{3}))\cdot up_{proj}(h_{3})]^{T}.$ (15)
According to Eq. 5-11, the gradient of $W_{n}$ is closely related to the value of $h_{1}$. Thus, when Gaussion noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$ as $\tilde{h}_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations. In other words, injecting noise into the forward propagation process can protect the hidden states and its gradients transmitted during the training of the splitting LLMs.
3.3 Parallel Training
(a) Client-batch parallel training.
(b) Server-hierachical parallel training.
Figure 5: Multi-client Parallel Training of FL-LLaMA. In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation. Then, the server splits hidden state $h_{B}$ and sends them back to different clients. In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$. After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
In multi-clients settings, the FL-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism and server-hierarchical parallelism. Each of them optimizes for different system configurations and resource availability. In sequential training, the server interacts with one client at a time in a round-robin manner, which is more suitable for scenarios with limited server capacity. In client-batch parallelism, the server waits for data from each client and then concatenate them in a single batch, which is appropriate for single machine with multiple GPUs server scenarios. In server-hierarchical parallelism, each client is assigned a separate sub-server for interaction, allowing multiple clients to train in true parallel, which is suitable for multi-machine, multi-GPU server environments.
3.3.1 Sequential Training
The client begins by passing hidden states through its input blocks $Blocks_{A}$ and forwards the result to the server. The server processes the input through $Blocks_{B}$ and returns the output states. The client then computes the loss with $Blocks_{C}$, sends the gradients to the server for model updates, and receives the updated gradients to adjust its input blocks. This process continues until all clients have completed their respective training phases for the current round. Although this method is inherently slower due to its serialized nature, it significantly reduces the computational burden on the server, making it a suitable choice for scenarios where server-side resources are limited.
3.3.2 Client-Batch Parallelism
[…]
3.3.3 Server-hierarchical Parallelism
In server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Fig. 5(b). These instances run in parallel, allowing each client to perform local training simultaneously rather than waiting for others to complete. Typically after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers. This is usually done via weighted averaging of both server-side and client-side model parameters. This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
Dataset Task Cloze Question Answers ReCoRD Question answering [passage p] [cloze question q] Answer candidates COPA Causal reasoning “[choice c1]” or“[choice c2]”? [premise p], so [M]. c1/c2 WSC Coreference resolution [sentence s] The pronoun ‘*p*’ refers to [M]. Noun n RTE Textual entailment “[hypothesis h]”? [M] “[premise p]” “yes”/“no” BoolQ Question answering [passage p]. Question: q? Answer:[M]. “yes” / “no” WiC Word sense disambiguation “[sentence s1]”/“[sentence s2]”Similar sense of [word w]? [M]. “yes”/“no” CB Textual entailment “[hypothesis h]”? [M], “[premise p]” “yes”/“no”/“maybe” MultiRC Question answering [passage p]. Question: q? Is it [answer a]? [M]. “yes”/“no”
Table 1: Cloze questions and answers for the 8 SuperGLUE tasks
3.4 Inference Accelaration
The inference time of LLMs are critically important for both enterprises and end-users. Low-latency inference minimizes user wait times, allowing for efficient and fast handling of complex tasks. Distributed architectures partition tensor operations, traditionally executed within GPUs and directly interacting with high-bandwidth memory (HBM), into separate computational units across multiple physical devices. Consequently, intermediate data must be transferred over networks, posing substantial latency challenges for the inference of LLMs. To enhance the efficiency and usability of distributed LLMs, our proposed FL-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
3.4.1 Attention Mask Compression
FL-LLaMA transmits only the minimal information necessary for effective model training and inference. One significant improvement arises from reducing redundancy in attention mask transmissions. In masked self-attention mechanisms, the attention mask serves two primary purposes: (1) preventing decoders from prematurely accessing subsequent tokens during prediction, and (2) ensuring that padding tokens within input data do not interfere with attention computations involving valid tokens. Notably, the generation of attention masks across transformer blocks remains consistent throughout each training iteration and depends solely on dataset padding. Consequently, rather than transmitting the entire attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$, which incurs substantial bandwidth latency, the FL-LLaMA framework only sends essential metadata ($seq\_len$, $pad\_len$) indicating the locations of padding tokens versus valid tokens from the client to the server, thus significantly decreasing communication overhead.
3.4.2 Collaborative KV Cache Mechanism
Since the model autoregressively generates the next token based on the tokens already generated, the speed of generating new tokens slows down as the input sequence length becomes excessively long (e.g., greater than 1000), due to the increasing number of K and V calculations. The application of key-value cache (KV cache) techniques in distributed inference further enhances inference speed. Generative LLMs produce outputs autoregressively, meaning that each generated token becomes part of the input sequence used to generate subsequent tokens. As user queries grow longer, repeatedly transmitting the complete context—comprising the initial query and previously generated tokens—from client to server can drastically increase inference latency. To mitigate this issue, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers. Since these states remain unchanged throughout token generation, newly generated tokens reuse cached states without requiring recomputation. Building upon this insight, FL-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step. This approach effectively minimizes data transmission, substantially improving inference efficiency.
3.5 Adaptive Partition
Ege et al.[49] and Dario et al.[50] have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations. Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks in accordance with the demands of downstream tasks and available computational resources, thereby achieving an optimal balance between performance and efficiency. Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with Blocks_A=1,2,3) and the last blocks (with Blocks_C=1,2,3) of the model in Sec. 4.2.6. The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients. Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
4 Experiments
In order to demonstrate the performance of LLaMA within the federated split framework (FL-LLaMA), we conduct experiments using the same benchmarks as those used in LLaMA2-7B model [53]. Developed by Meta AI, the LLaMA2-7B model serves as a general-purpose language model with strong performance in natural language processing tasks, particularly when fine-tuned for specific domains. We employ this architecture to investigate whether FL-LLaMA can maintain comparable task performance to conventional centralized training, while addressing high delays and data privacy concerns through decentralized learning.
Table 2: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FL-LLaMA and baselines.
[…]
4.1 Experimental Settings
In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FL-LLaMA.
4.1.1 Dataset
To evaluate model generalization across diverse NLP tasks, our experiments incorporates three distinct tasks:
  • 1.
    The SuperGLUE benchmark [58] is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, which span a broad range of linguistic challenges, from common sense reasoning to fine-grained semantic interpretation. Table 1 illustrates examples of cloze-style questions and corresponding answers derived from each task. (1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages. (2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise. (3)WSC focuses on pronoun disambiguation requiring deep contextual understanding. (4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another. (5)BoolQ involves yes/no questions answered using evidence from context passages. (6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences. (7)CB deals with entailment in complex sentences with nuanced reasoning. (8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information. To unify these tasks under a single modeling framework, we adopt the formulation used in GLM [56], representing each instance as a blank infilling problem. Given a labeled pair $(x,y)$, the input $x$ is converted into a masked cloze question $q(x)$ using a placeholder token $[M]$, while the output $y$ serves as the target answer $a(y)$.
  • 2.
    The CoQA dataset [59] is designed for the task of conversational question answering(QA), focusing on evaluating contextual coherence in multi-turn dialogues. It includes over 8,000 dialogues and more than 127,000 questions drawn from seven diverse domains. Answers are provided in free-form text, with nearly half of the questions requiring coreference resolution (e.g., pronoun disambiguation) and pragmatic reasoning. This setup closely mirrors real-world conversational settings, aiming to assess a model’s ability to understand text passages and respond to a series of interrelated questions.
  • 3.
    The XSum dataset [60] is designed for the task of abstractive summarization, specifically evaluating a model’s ability to perform extreme compression of salient information from news articles. It consists of BBC news articles paired with manually written summaries, covering a broad range of topics. A key challenge of XSum lies in its requirement for the model to generate a single-sentence summary for each article, placing significant demands on the model’s ability to abstract and condense information effectively.
4.1.2 Metrics
Since NLU tasks are reformulated as blank infilling problems, model performance is assessed based on the probability assigned to the ground-truth answer $a(y)$. For datasets such as RTE, BoolQ, WiC, CB, and MultiRC where the correct answer is often a single token, the evaluation score is derived from the logit corresponding to that token. Specifically, we define the probability as follows:
$p(y|x)=\frac{p(a(y)|q(x))}{\sum_{y^{{}^{\prime}}\in Y}p(a(y^{{}^{\prime}})|q(x))},$
where $Y$ represents the set of possible ground-truth labels.
In contrast, datasets like ReCoRD, COPA, and WSC require multi-token answers. In these cases, we evaluate model performance by summing the log-probabilities of each token in the predicted answer, computed as:
$s(y|x)=\sum_{t=1}^{|L_{y}|}\text{log}\text{P}(y_{t}|y_{1}...y_{t-1},x;\theta).$
For conversational QA tasks, we utilize two metrics: Exact Match (EM) and F1 score to compare the model performance, which evaluate the accuracy of predicted answers against the ground truth [61, 62]. Exact Match (EM) measures the percentage of predictions that match the ground truth answers exactly, accounting for variations in punctuation and case. The F1 score is the harmonic mean of precision and recall, providing a more nuanced view of performance by rewarding partial overlaps between the predicted and reference answers.
For summarization tasks, we adopt standard ROUGE metrics [63, 64, 65], i.e., ROUGE-1 and ROUGE-2, which are widely used to measure content overlap between generated summaries and reference texts. ROUGE-1 evaluates the unigram overlap, while ROUGE-2 measures the bigram overlap, providing insights into both content coverage and fluency.
4.1.3 Baselines
[…]
Table 3: Comparison of training time between different training strategies on ReCoRD dataset.
Strategy Sequential training Client-batch training Server-hierachical training   Centralized
num. of clients 2 2 4 8 2 4 8 -
time(s) 689.2$\pm$5.7 630.0$\pm$1.8 542.2$\pm$3.5 488.6$\pm$0.4 348.2$\pm$3.8 173.53$\pm$3.0 86.4$\pm$0.6 381.9$\pm$2.4
4.1.4 Parameter Settings
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation of the FL-LLaMA is based on the Flower [67] federated learning framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters.
4.2 Experimental Results
4.2.1 FL-LLaMA Performance
In order to validate the effectiveness of FL-LLaMA, we compared the performance of FL-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 2. Experimental results show that FL-LLaMA and centralized LLaMA2-7B exhibit no significant performance differences, demonstrating the feasibility of the framework. Not only does FL-LLaMA perform comparably to centralized LLaMA2-7B, but the GPU memory consumption on the client side is also significantly reduced compared to centralized LLaMA2-7B. The comparison of their GPU memory usage is shown in Fig. 6. While centralized LLaMA2-7B occupies 27.9G of GPU memory, the FL-LLaMA client requires only 6G, representing a reduction of up to 82% in memory requirements. This reduction enables enterprises or departments that possess large amounts of high-quality private datasets but have limited computational resources to participate in federated training of LLMs.
Figure 6: Computation load on client side of centralized LLaMA2-7B and FL-LLaMA on ReCoRD dataset.
4.2.2 Parallel Training
For communication cost analysis, we randomly selected 1,000 samples from the ReCoRD dataset. In our experimental setup, all training used a batch size of 1 and adopted FP16 training to improve computational efficiency. For LoRA, the decomposition rank $r$ is set to 8, and the scaling factor $\alpha$ is configured to 16, focusing specifically on the attention modules during fine-tuning. Additionally, the model is partitioned with Blocks_A=1 and Blocks_C=1. Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time, as summarized in Tab. 3. The results show that sequential training takes the longest time, followed by client-batch training, and server hierarchical takes the shortest time. However, when the server can accommodate a larger batch size, increasing the number of clients in client-batch training effectively increases the batch size and thus reduces overall training time.
Figure 7: Avg magnitude of hidden state $h_{A}$ (Blocks_C =1) on MultiRC dataset.
(a) Loss Curves for Noise Scales
(Blocks_A=1, Blocks_C=1)
(b) Loss Curves for Noise Scales
(Blocks_A=2, Blocks_C=1)
(c) Loss Curves for Noise Scales
(Blocks_A=3, Blocks_C=1)
Figure 8: FL-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
4.2.3 Effects of Noise Scales
Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how varying levels of noise affect model performance. The results demonstrate that introducing appropriate levels of noise leads to only a slight decrease in the performance of distributed LLMs, while significantly enhancing their security.
The magnitude of noise injection significantly influences the stability and convergence speed of model training. Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance. To quantify this effect, we first fine-tuned the FL-LLaMA model on the MultiRC [68] with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within hidden state $h_{A}$. According to Fig. 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1). The average magnitude of hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$. Based on this observation, we subsequently injected noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We used LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2.
The corresponding experimental results are presented in Fig. 8(a)-8(c) and Tab. 4. Fig. 8(a)-8(c) demonstrate that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ are held constant, the convergence speed of FL-LLaMA decreases as the magnitude of injected noise increases. When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline. A horizontal comparison reveals that, under the same noise level, increasing the number of blcoks in $Blocks_{A}$ consistently accelerates convergence.
Tab. 4 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores. Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness. This can be explained by the average magnitude shown in Fig. 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR). This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
Table 4: Different noise scales and blocks number of $Blocks_{A}$ analysis(F1a / EM) on MultiRC dataset.
Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3
0.0 77.48/40.82 77.82/39.24 77.89/41.34
0.02 77.89/38.09 77.87/39.66 78.61/42.39
0.05 74.57/29.17 77.07/38.41 78.95/41.03
0.1 59.11/8.18 72.45/23.71 75.68/34.10
4.2.4 Model Reverse Attack
To investigate the impact of split point selection and noise scale on the security of FL-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5). The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5. Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2. FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57. Experiment results of FL-LLaMA is shown in Tab. 5. From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores. However, under the FL-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero. Tab. 6 reports the downstream performance of $client_{0}$ after multi-client training, showing that secret attack allows training to proceed as usual without disrupting the federated process. These findings demonstrate that FL-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
Table 5: Security analysis (Rouge-2 F1 / BlEU-4) of FL-LLaMA on Record dataset.
Noise scale Blocks_A=1 Blocks_A=2 Blocks_A=3
0.0 0.051/0.003 0.056/0.003 0.069/0.016
0.02 0.050/0.002 0.052/0.003 0.084/0.044
0.05 0.050/0.003 0.051/0.002 0.067/0.004
Table 6: Performance (EM) under different noise scales and blocks number of $Blocks_{A}$ on Record dataset.
Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3
0.0 77.4 79.8 78.9
0.02 79.1 79.9 79.6
0.05 79.6 79.4 79.2
4.2.5 Efficiency Optimization
Figure 9: FL-LLaMA: Inference Speed Comparison (w/ vs. w/o Attention Mask).
Fig. 9 shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training. The bar chart represents the average time spent per token in FL-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second by FL-LLaMA. It can be observed from the figure that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
Figure 10: FL-LLaMA: Inference Speed Comparison(w/ vs. w/o KV Cache).
Fig. 10 illustrates the effect of the KV cache technique on improving inference speed for different query lengths in distributed inference. The experimental results indicate that without the KV cache, the inference speed of FL-LLaMA significantly decreases as the query length increases, primarily due to the continuously increasing number of K and V vectors computed at each step in the self-attention mechanism. In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing the previously generated K and V vectors, thereby effectively avoiding redundant computations. When the query length reaches 4000, the inference speed of FL-LLaMA with KV cache is 8.2 times that without it, and as the query length increases further, this advantage will continue to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
4.2.6 Adaptive Splitting
The choice of model partitioning has a significant impact on the computational burden at the client side. To evaluate whether FL-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3). This setting reflects realistic constraints where clients possess limited computational resources. All models are trained for 8,000 steps using FP16 precision. LoRA is applied with a rank of 16 and a scaling factor of $\alpha=32$. The training used a batch size of 2 and a learning rate of 2e-5.
Table 7: Different model partition analysis(F1a/EM) on MultiRC dataset. No Noise is added to the hidden states.
Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3
1 80.34/44.60 80.63/46.59 80.90/47.11
2 80.41/45.86 80.62/45.96 79.94/45.02
3 81.17/47.22 80.74/46.59 80.20/45.33
Tab. 7 presents F1a/EM scores under various partitioning schemes. The highest F1a/EM scores (81.17/47.22) emerge when Blocks_A=3 and Blocks_C = 1, indicating that allocating more early blocks to the client side can slightly enhance task performance. Overall, the performance differences across partitions remain small, suggesting that FL-LLaMA is relatively insensitive to how the model is split. From a practical standpoint, these findings confirm that clients with constrained resources can flexibly adjust how many blocks they store without substantially degrading accuracy. In scenarios where client-side memory or compute capacities are limited, reducing Blocks_A and Blocks_C remains viable, given that FL-LLaMA maintains robust performance across a range of partitions.
Building upon partitioning experiment, we introduced controlled noise into the hidden states produced by $Blocks_{A}$. Specifically, Gaussian noise with a scale of 0.02 is added during forward propagation before sending the hidden states to the server. Tab. 8 shows that while there is a slight drop compared to the noise-free scenario, the overall performance remains robust, indicating that moderate perturbations in the hidden states do not severely impact convergence or accuracy. Futher, the experimental results show that the higher the number of $Blocks_{A}$, the smaller the performance degradation of FL-LLaMA on downstream tasks. This indicates that the resistance to noise is higher, which verifies that the signal-to-noise ratio has a significant impact on the performance of the model.
Table 8: Different model partitions analysis(F1a/EM) on MultiRC dataset. Noise (scale=0.02) is added to the states $h_{A}$.
Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3
1 79.44/38.41 79.58/38.61 79.45/39.66
2 80.10/41.55 80.14/38.61 79.99/40.40
3 79.90/40.29 80.29/43.02 80.44/42.18
5 Conclusion
In this work, we proposed FL-LLaMA, a novel federated split learning framework tailored for large language models. By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FL-LLaMA addresses the core limitations of traditional federated split learning approaches. Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FL-LLaMA demonstrates significant advances in security, efficiency, and adaptability. This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs. In future work, we aim to support more basement models and deploy FL-LLaMA framework in scenarios where data privacy is a critical concern.

### v2

@@
-
+FedSEA-LLaMA: A Secure, Efficient and Adaptive Federated Splitting Framework for Large Language Models
 Abstract
-Private data is typically larger and of higher quality than public data, offering great potential to improve LLM.
-However, its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments.
-To address this, the transformer-based split learning model has emerged, offloading most model parameters to the server while retaining only the embedding and output layers on clients to ensure privacy.
-However, it still faces significant challenges in security, efficiency, and adaptability: 1) embedding gradients are vulnerable to attacks, leading to reverse engineering of private data; 2) the autoregressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) fixed partition points lack adaptability to downstream tasks.
-In this paper, we introduce FL-LLaMA, a secure, efficient, and adaptive federated split framework based on LLaMA2.
-First, we place some input and output blocks on the local client and inject Gaussian noise into forward-pass hidden states, enabling secure end-to-end propagation.
-Second, we employ client-batch and server-hierarchical strategies to achieve parallel training, along with attention-mask compression and KV cache mechanisms to accelerate inference, reducing communication costs effectively.
-Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements and hardware limitations.
-Experiments on NLU, summarization and conversational QA tasks show that FL-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 2x train speedups and 8× inference speedups.
-Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FL-LLaMA in security and adaptability.
-keywords:
-Federated Learning , Large Language Model , Privacy Protection , Communication Efficiency , Split Learning
-PACS:
-07.05.Mh , 84.35.+i
-MSC:
-[2010] 68Q85 , 68W15 , 68T05 , 68T50 , 68W10
-journal: Artificial Intelligence
+Private data holds promise for improving LLMs due to its high quality, but its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments.
+To address this, the transformer-based federated split models are proposed, which offload most model parameters to the server (or distributed clients) while retaining only a small portion on the client to ensure data privacy.
+Despite this design, they still face three challenges: 1) Peer-to-peer key encryption struggles to secure transmitted vectors effectively; 2) The auto-regressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) Fixed partition points lack adaptability to downstream tasks.
+In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated splitting framework based on LLaMA2.
+First, we inject Gaussian noise into forward-pass hidden states to enable secure end-to-end vector transmission.
+Second, we employ attention-mask compression and KV cache collaboration to reduce communication costs, accelerating training and inference.
+Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements.
+Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups of training and inference.
+Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FedSEA-LLaMA in security and adaptability.
+Introduction
+While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (Kaplan et al.
+2020), the true potential of these models lies in their application to high-quality private data.
+Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (Yang et al.
+2024).
+Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (Ye et al.
+2024).
+Therefore, effectively incorporating private data will further advance the performance of LLMs.
 […]
-[…]
-{graphicalabstract}[Uncaptioned image]
-[…]
-{highlights}
-We introduce the federated splitting LLM, FL-LLaMA.
-FL-LLaMA holds some input and output blocks on the local client and injects Gaussian noise into forward-pass hidden states, enabling secure end-to-end propagation.
-We employ client-batch and server-hierarchical strategies to achieve parallel training, along with attention-mask compression and KV cache mechanisms to accelerate training and inference.
-FL-LLaMA allows users to dynamically adjust the partition points for input/output blocks based on specific task requirements and hardware limitations.
-1 Introduction
-The success of Large Language Models (LLMs) can be attributed to their ability to leverage large amounts of public data and intricate parameter networks [1].
-By training on extensive corpora, LLMs can capture statistical patterns in language, semantic structures, and subtle contextual relationships.
-However, private data holds a significant advantage over public datasets in terms of both quality and utibility.
-This is because private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies [2].
-Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios [3].
-Therefore, effectively incorporating private data will become one of the critical factors in further advancing the performance of LLMs.
-Figure 1: Model architecture of FedBert and FL-LLaMA.
 The distributed nature of private data across various silos and the significant computational demands of LLMs present substantial challenges for their deployment in federated environments.
-Private data is often isolated in local silos, such as mobile devices [4], enterprises servers [5] or medical institutions [6, 7], as centralized storage and processing of such data may raise privacy concerns and regulatory risks [8].
-Federated Learning (FL) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation, but traditional FL frameworks, such as FedAvg [9] and FedProx [10], are infeasible for LLMs training scenarios.
-LLMs typically have billions of parameters, and their training process requires extensive computational resources, placing substantial demands on device memory and network bandwidth.
-This makes training the full model on ordinary client devices practically infeasible.
-Additionally, the communication overhead and heterogeneity inherent in LLMs training become even more pronounced within federated environments, making it challenging for models to efficiently converge or maintain generalization capabilities.
-As a result, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open and cutting-edge research topic.
-As an alternative method, transformer-based split learning methods, named FedBERT [11], have been proposed to adapt LLMs, offloading most model parameters reside to the server while retaining only the embedding and output layers on client to ensure privacy.
-As shown in Fig.
-1(a), FedBERT divides Transformers into three parts, i.e., embedding layers and output layers on client, while main Transformer layers on server.
-The forward propagation(FP) path is client-to-server-to-client.
-After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client.
-The core idea of FedBERT is to divide the calculation into two parts: the client devices is responsible for initial feature extraction from raw data and calculate final loss, while the server handles deeper inference and learning.
-Although FedBERT provides a potentially viable solution for resource-constrained clients scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the embedding layer are transmitted, which can be exploited to reconstruct original input data through beam search [12] or reverse engineering [13].
+Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprises servers (Huang et al.
+2024) or medical institutions (Qi et al.
+2024; Yuan et al.
+2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al.
+2024).
+Federated Learning (FL) (Stremmel and Singh 2021; Li et al.
+2020) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation.
+They often leverage Adapters (Houlsby et al.
+2019) or LoRA (Hu et al.
+2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to achieve fine-tuning LLMs.
+However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible.
+Therefore, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open research problem.
+To address this, transformer-based federated splitting models (Zheng et al.
+2024; Su et al.
+2024) have been proposed to adapt LLMs fine-tuning to the resource-constrained scenarios.
+Titanic (Su et al.
+2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training.
+FL-GLM (Zheng et al.
+2024) offloads most model parameters to resource-ample server while retaining the input and output layer on client to ensure data privacy.
+As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on client, while the middle blocks on the server.
+The forward propagation(FP) path is client-to-server-to-client with peer-to-peer key encryption.
+After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client with the same encryption.
+Although FL-GLM provides a potentially viable solution for clients resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al.
+2023).
 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens.
 In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage.
-3) Lack of Adaptability: Fixed partition points lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
-In this paper, we introduce FL-LLaMA, a secure, efficient, and adaptive federated split framework based on LLaMA2, as shown in Fig.
-1(b).
-To Enable Secure Propagation, we partition the transformers of LLaMA2 into three parts: input and output blocks on clients, while the remaining large parameters on server.
-During training, clients perform forward propagation to generate hidden states, add Gaussian noise for them, and send these noised states to the server for further processing.
-Through both gradients analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages.
-To Reduce Communication Cost, we employ client-batch and server-hierarchical strategies to achieve parallel training: the former batches data from different clients, while the latter leverages concurrent sub-server to enhance training efficiency.
-Durning inference, we introduce the attention-mask compressio and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of autoregressive inference.
-By synchronously maintaining cache states on both client and server, the originally cross-device transmition is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B).
+3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
+In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated split framework based on LLaMA2, as shown in Figure 1b.
+To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks, and send these noised states to the server for further processing.
+Through both gradient analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages.
+To Reduce Communication Cost, we introduce the attention-mask compression and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of auto-regressive inference.
+By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B).
 To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources.
-This adaptability enables optimal balance between performance and efficiency.
-Experiments on NLU, summarization, and conversational QA tasks show that FL-LLaMA maintains performance comparable to centralized LLaMA2, and achieving up to 2x training and 8× inference speedups.
-Through inference acceleration, the average token generation time of FL-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the length of context.
-[^2] Our code is available at: https://github.com/TAP-LLM/SplitFedLLM.
+This adaptability enables an optimal balance between performance and efficiency.
+Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 8× inference speedups.
+Through inference acceleration, the average token generation time of FedSEA-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the growing length of context.
+Additional analysis of privacy attacks and various partition points further highlights FedSEA-LLaMA’s effectiveness in terms of security and adaptability.
+[^1] Our code is available after the submission is accepted
 The innovations in this paper are as follows:
-  • 1.
-    We design a federated split framework tailored for resource-constrained clients scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
-  • 2.
-    We propose client-batch and server-hierarchical optimizations to enhance split learning efficiency, and use collaborative KV cache with attention mask compression to reduce autoregressive inference communication.
-  • 3.
-    We enables dynamic partition points based on downstream tasks and device capabilities, optimizing the trade-off between computational efficiency and model performance.
-  • 4.
-    Experimental results on NLU, summarization and conversation QA tasks show that our FL-LLaMA achieves performance comparable to centralized LLaMA2, and achieve up to 2x train speedups and 8× inference speedups.
-This paper is an extension of the work presented in  [14].
-There are five major differences from the original version: (i) the base model and fine-tuning method have been updated, from p-tuning v2 on ChatGLM in the original version to LoRA on LLaMA2, demonstrating the generalizability of our federated split framework; (ii) Gaussian noise is introduced during the forward pass of split learning to further enhance the security of the federated split LLMs; (iii) the use of attention-mask compression and collaborative KV cache mechanism significantly improves inference efficiency in the federated split setting; (iv) dynamic partitioning is supported, allowing users to customize split points via configuration files, thus improving adaptability; (v) the experimental evaluation has been extended by assessing FL-LLaMA on new summarization and conversational QA tasks, showcasing its versatility.
-The rest of this paper is organized as follows.
-Section 2 reviews related work, and Section 3 presents the details of the federated split framework.
-Section 4 reports the experimental evaluation.
-Finally, Section 5 concludes the paper.
-2 Related work
-2.1 Fedrated Learning for LLMs
-Federated Learning (FL) has emerged as a promising paradigm for training language model in a decentralized manner while preserving user privacy and ensuring data security.
-FedAvg [15] is one of the earliest and most widely adopted optimization algorithms, which has been successfully applied in multiple language modeling tasks [16, 17, 9].
-In this approach, each client locally trains a model on its private data and sends the updates to a central server, which performs a weighted aggregation to produce a new global model.
-To reduce the number of local training rounds and accelerate convergence, Stremmel et al.
-[9] propose initializing the global model with a pretrained language model before applying FedAvg, which has proven effective in speeding up training.
-Ji et al.
-[18] introduce FedAtt, which incorporates a hierarchical soft attention mechanism during aggregation to adaptively weighs each client’s model parameters based on their contribution, improving both generalization and local adaptation.
-In terms of LLMs, traditional FL frameworks face growing challenges due to increased communication and computational costs.
-To address the issue of massive model sizes, some federated LLMs framework  [19, 20, 3] has been proposed to leverage instructions tuning, Adapters and LoRA for local fine-tuning, which require updating only a small subset of parameters and significantly reducing communication costs while maintaining performance.
-But these methods still require the transmission of parameter quantities far exceeding the system’s carrying capacity.
-To further minimize communication overhead, gradient compression [21, 22] and quantization [23] mechanisms have been explored, which effectively reduce the volume of data exchanged, making them well-suited for federated fine-tuning of LLMs.
-Liu et al.
-[24] propose structural pruning approaches that update only selected parameter layers of LLMs based on Fisher Information.
-By narrowing the update scope, these methods effectively reduce the size of transmitted models without sacrificing overall performance.
-Besides knowledge distillation methods between large and small language models [25, 26, 27] are proposed to avoid the high communication cost of transferring LLMs parameters directly.
-It utilize the smaller proxy models acting as intermediaries, facilitating knowledge transfer between clients and the server with substantially lower communication.
-Additionally, client heterogeneity remains a central challenge in FL, due to variations in computational capacity, data distributions, and task types across clients [28].
-Bai et al.
-[29] and Wang et al.
-[30] introduce LoRA-based stacking strategies to balance the differences in client capabilities.
-They help ensure that models from clients with varying ranks can still be aggregated effectively, improving training stability and fairness.
-While recent works of federated LLMs explores efficient fine-tuning and parameter reduction methods, resource-constrained clients still struggle with independent fine-tuning, especially with added compression overhead.
-Therefore, split learning has emerged as a viable alternative for such scenarios.
-2.2 Split Learning for LLMs
-Split learning [31] is gaining traction as a decentralized and collaborative training paradigm tailored for environments with limited computational resources [32, 33, 34, 35], particularly suited for scenarios involving mobile platforms or lightweight clients that lack access to GPUs.
-The core idea involves partitioning a deep neural network into separate sub-network, with each sub-network assigned to a different device for distributed processing.
-To address the training of sequential data in LLMs, FedBERT [11] introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices.
-In this approach, the BERT model is divided into three core components: the Embedding layer, the Transformer layer, and the Output layer.
-This separation allows the heavy Transformer layer to be trained on a centralized, high-performance server, while the lighter Embedding and Output layers are handled locally by client devices.
+  • We design a federated split framework tailored for resource-limited scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
+  • We leverage collaborative KV caching and dynamic attention mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
+  • Experimental results on different three tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency and adaptability of our framework.
+Related Works
+[…]
+Federated Learning for LLMs
+Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (McMahan et al.
+2017; Stremmel and Singh 2021; Ji et al.
+2019).
+Recent federated LLM frameworks (Fan et al.
+2023; Kuang et al.
+2024; Ye et al.
+2024) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead.
+However, transmission demands often still exceed system limits.
+To mitigate this, gradient compression (Wu et al.
+2024b; Shu et al.
+2024) and quantization (JianHao et al.
+2024) techniques have been applied.
+Structural pruning based on Fisher Information (Liu et al.
+2024) further narrows updates to key layers, reducing model size without performance loss.
+Knowledge distillation (Wu et al.
+2024a; Fan et al.
+2024, 2025) offers another solution, using smaller proxy models to transfer knowledge with minimal communication.
+Client heterogeneity remains a major challenge (Zhang et al.
+2024), prompting methods like LoRA-based stacking (Bai et al.
+2024; Wang et al.
+2024) to align models from diverse clients and improve fairness.
+Despite advances in efficient fine-tuning and parameter reduction for federated LLMs, resource-constrained clients still face challenges with independent fine-tuning, particularly under compression overhead.
+Split learning (Thapa et al.
+2022) offers a promising alternative in such settings.
+Split Learning for LLMs
+Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al.
+2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients.
+It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning.
+To address the training of sequential data in LLMs, FedBERT (Tian et al.
+2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices.
 Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points and susceptibility to privacy threats like embedding gradient attacks.
 Recently, Su et al.
-[36] extend the scope of split learning by exploring scenarios with heterogeneous resource configurations, particularly considering a case where the server lacks sufficient computational capacity.
+(2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity.
 In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation.
-Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time for LLMs.
-Furthermore, the repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities.
-Our work diverges from these studies by targeting a scenario with a powerful server and resource-constrained clients holding private data.
-We prioritize enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
-[…]
-2.3 Attacks and Defenses
-In federated learning, although it allows multiple clients to jointly fine-tune models without revealing raw data, there remains a risk of embedding-based reconstruction attacks, LoRA-based gradient inversion attacks and split learning-based data reconstructing attacks.
-Attacks from embedding gradients have proven highly effective in reconstructing users’ private data.
-Gupta et al.
-[12] demonstrated that by identifying non-zero elements in the embedding gradients, an adversary can infer which tokens were used by the client, using beam search to reconstruct private data.
-To mitigate such attacks, they proposed freezing the embedding layer during training.
-Zhu et al.
-[37] briefly discussed defensive strategies such as injecting differential privacy noise [38, 39] or applying gradient clipping by zeroing out small gradients.
-Huang et al.
-[40] introduced a MixUp-based data augmentation method on the [CLS] token in BERT, while Yaldiz et al.
-[41] suggested applying cosine similarity checks on client-uploaded weights to detect and filter out malicious clients.
-But these defenses often degrade the model’s accuracy [42, 43].
-In addition to embedding-based reconstruction attacks, Federated LLMs are also vulnerable to gradient inversion attacks via LoRA, due to frequent parameter sharing across clients [44, 45].
-To address this, Liu et al.
-[46] propose DP-LoRA, which adds Gaussian noise to LoRA updates to ensure differential privacy [47] [39].
-However, the combination of heterogeneous data, DP noise, and LoRA’s sensitivity to hyperparameters can lead to instability and performance loss.
-To overcome this, Sun et al.
-[48] introduced FFA-LoRA, which initializes the non-zero entries of the LoRA $A$ matrix from a Gaussian distribution and keeps matrix $A$ frozen while only fine-tuning the zero-initialized LoRA $B$ matrix, thus preventing the injected noise under differential privacy from being amplified.
-However, exposure of intermediate activations and gradients in distributed training still enables potential data reconstruction, prompting increased focus on secure split learning methods.
-In split learning, Ege et al.
-[49] propose inferring labels by minimizing gradient differences between a pseudo-model and the actual model, achieving strong performance in classification tasks.
-However, this approach is impractical for LLMs due to the high computational cost of handling vocabularies exceeding 100,000 tokens, making real-time federated training infeasible.
-In image recognition, Dario et al.
-[50] show that a malicious server could reconstruct client data by exploiting hidden states via a Feature Space Hijacking Attack (FSHA).
-However, FSHA relies on access to a shadow dataset from the same domain and task—an unrealistic assumption in privacy-preserving settings, where servers lack access to such data.
-Building on prior attacks, we explore a more severe threat model where a malicious client colludes with an honest-but-curious server by sharing its data, enabling the server to train an attack model that reconstructs plaintext from other clients using hidden states.
-Importantly, the server does not alter model outputs or gradients.
-3 Framework
-In this section, we present a detailed description of the FL-LLaMA framework, as shown in Fig.
-2, which comprises five components: split learning, privacy preservation, parallel training, inference acceleration and adaptive partitioning.
-First, we adopt split learning to divide the LLaMA into three parts: the local client stores the first ${p}$ blocks and the last ${q}$ blocks, while the remaining ${k}$ blocks, comprising the majority of the model parameters, are hosted on the server.
+Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time.
+Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities.
+Zheng et al.
+(2024) propose FL-GLM to place the input and output blocks locally on client devices, while remaining primary model parameters are centralized on a server with ample computational resources.
+And then, they employ key-encryption during client-server communication to prevent reverse engineering attacks from peer clients.
+But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
+Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
+Model
+In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration and adaptive partitioning.
+Followed FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks and the last q{q} blocks, while the remaining k{k} blocks, comprising the majority of the model parameters, are hosted on the server.
 Next, we design a Gaussian noise injection mechanism applied to the forward-pass hidden states to ensure data privacy during transmission between the client and server.
-Third, we enhance the training efficiency of split learning through a client-batch and server-hierarchical parallelisms.
-Fourth, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms.
-Finally, the framework allows clients to adaptively adjust the split configuration based on their privacy preferences and available computational resources.
-3.1 Split Learning
-Suppossing that the LLMs comprises $N$ transformer blocks, the proposed FL-LLaMA framework partitions the LLMs into three segments: Part I contains the embedding layer and the first $p$ transformer blocks, denoted as $Blocks_{A}$; Part II includes the middle $k$ transformer blocks, denoted as $Blocks_{B}$; Part III consists of the last $q$ transformer blocks along with the output layer, denoted as $Blocks_{C}$, where $p+k+q=N$.
-For generative tasks, the model predicts the next token autoregressively, with data labels closely linked to the input text sequences.
-To safeguard client data privacy, the FL-LLaMA framework places Part I $Blocks_{A}$ and Part III $Blocks_{C}$ on the client side, while Part II $Blocks_{B}$ resides on the server side, thereby preventing direct exposure of client data.
-Such an architecture optimizes GPU memory allocation, as the server handles most of the model’s GPU memory burden, significantly reducing memory pressure on the client.
-The architecture diagram of the proposed model is shown on the left of Fig.
-2.
-Taking a single-client distributed training scenario as an example, at the beginning of training, the client utilizes local data to perform forward propagation through $Blocks_{A}$, resulting in hidden states $h_{A}$:
-$h_{A}=f_{1\sim p}\circ f_{\text{embed}}(\mathbf{x}),$ (1)
-where $f_{\text{embed}}(\cdot)$ denotes the embedding layer of the LLMs, and $f_{1\sim p}(\cdot)$ represents the sequential forward propagation through transformer block 1 to block $p$ at the client side.
-After obtaining $h_{A}$, the client transmits it together with the position index and the attention mask matrix $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ to the server, where $b$ is the batch size and $seq\_len$ is the maximum context length in LLMs.
-Subsequently, $Blocks_{B}$ on the server side takes the received hidden states $h_{A}$ and attention mask $A_{mask}$ as inputs, performs computationally intensive tasks, and returns the resulting hidden states $h_{B}$ back to the client:
-$h_{B}=f_{p+1\sim p+k}(h_{A},A_{mask}),$ (2)
-where $f_{p+1\sim p+k}(\cdot)$ denotes sequential forward propagation through transformer blocks $p+1$ to $p+k$ on the server side.
-Finally, $Blocks_{C}$ on the client-side performs computations to obtain the logits $\mathbf{z}$, computes the loss with local data, and initiates gradient backpropagation.
-These steps are defined by the following equations:
-$\mathbf{z}=f_{p+k+1\sim N}(h_{B},A_{mask}),$ (3)
-$\mathcal{L}=\mathcal{F}_{loss}(f_{out}(\mathbf{z}),\mathbf{y}),$ (4)
-where $f_{p+k+1\sim N}(\cdot)$ denotes sequential forward propagation through transformer blocks $p+k+1$ to $N$ on the client side, $f_{out}(\cdot)$ is the output layer of the model, $\mathbf{y}$ is the targeting sentences and $\mathcal{F}_{loss}(\cdot)$ denotes the cross-entropy loss function.
-After computing the loss, the client performs gradient backpropagation locally to update $Blocks_{C}$, then transmits the gradient of the hidden states $h_{B}$ back to the server.
-The server executes gradient backpropagation on $Blocks_{B}$, updates its parameters, and subsequently transmits the gradients of hidden states $h_{A}$ back to the client.
-Lastly, $Blocks_{A}$ performs gradient backpropagation, updates the model and embedding parameters, and completes a sequential training round.
-3.2 Privacy Preservation
+Third, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms.
+Privacy Preservation
 In distributed training of LLMs, the split learning paradigm partitions the entire model into multiple sub-models deployed across different physical nodes.
@@
 […]
-3.2.1 Threat Model Defination
-We design an extreme attack scenario, inspired by Dario et al.
-[50], to evaluate the robustness of FL-LLaMA.
-Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Fig.
-3, where a malicious client colludes with an honest-but-curious server.
+Threat Model Defination
+We design an extreme attack scenario, inspired by Pasquini, Ateniese, and Bernaschi (2021), to evaluate the robustness of FedSEA-LLaMA.
+Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server.
 The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally.
@@
 Concretely, we consider an adversary that has full knowledge of the client-side model architecture and the number of layers deployed on the client, but does not have access to the actual model parameters.
-Thus, the honest-but-curious server locally initializes a decoder with the same architecture and depth as the client’s model, same with LLaMA2 configuration, and uses it as the attack model.
-Without loss of generality, we consider a two-client FL-LLaMA setting, where the server receives hidden states from both $client_{0}$ and $client_{1}$.
+Thus, the honest-but-curious server locally initializes a decoder with the same architecture and depth as the client’s model and uses it as the attack model.
+Without loss of generality, we consider a two-client FedSEA-LLaMA setting, where the server receives hidden states from both $client_{0}$ and $client_{1}$.
 During each training step, the server first performs a normal forward pass through $Blocks_{B}$ to obtain the intermediate hidden states $h_{B}$.
@@
 Afterward, $h_{B}$ is split and sent back to the respective clients, allowing training to proceed as usual without disrupting the federated process.
-Figure 4: Inject noise to hidden state $h_{1}$. Noise added to $h_{1}$ eventually protects the gradients of parameters.
-3.2.2 Defense with Gaussion Noise Injection
-We explore two defense strategies to protect the transmission in FL-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states [^3] It is noteworthy that we chose not to use the standard DPSGD [51], which adds noise after computing all gradients, because FL-LLaMA requires noise protection during both forward and backward passes, unlike DPSGD’s focus on post-backpropagation gradients..
+Defense with Gaussion Noise Injection
+We explore two defense strategies to protect the transmission in FedSEA-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states.
 The first strategy introduces random Gaussian noise into the initial gradient during backpropagation (i.e., the gradient $\nabla h_{B}$), while the second applies Gaussian noise to the client’s first forward-pass output (i.e., the hidden state $h_{A}$).
@@
 In contrast, perturbing the hidden state $h_{A}$ during the forward pass yields more stable training.
-Therefore, FL-LLaMA adopts the latter to enhance robustness while preserving privacy.
+Therefore, FedSEA-LLaMA adopts the latter to enhance robustness while preserving privacy.
 Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation.
-Fig.
-4 illustrates the schematic of noise injection during the forward propagation of LLMs.
+Figure 3b illustrates the schematic of noise injection during the forward propagation of LLMs.
 Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer and $W_{n+1}$ for the weight of output linear layer.
@@
 $\hat{y}$ $=\text{softmax}(h_{5}),$ (6)
-$\mathcal{L}$ $=CrossEntropy(y,\hat{y}),$ (11)
+[…]
 where $\text{RMSNorm}(*)$ is the layernorm function, $\mathcal{A}$ is the attention layer, $\mathcal{M}$ is the MLP layer.
-During backpropagation, according to [52], the gradient of $W_{n}$ is:
+During backpropagation, according to (Clark 2017), the gradient of $W_{n}$ is:
 $\frac{\partial L}{\partial W_{n}}$ $=\frac{\partial L}{\partial h_{5}}\cdot\frac{\partial h_{5}}{\partial h_{4}}\cdot\frac{\partial h_{4}}{\partial W_{n}},$ (8)
@@
 According to Eq.
-5-11, the gradient of $W_{n}$ is closely related to the value of $h_{1}$.
-Thus, when Gaussion noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$ as $\tilde{h}_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations.
-In other words, injecting noise into the forward propagation process can protect the hidden states and its gradients transmitted during the training of the splitting LLMs.
-3.3 Parallel Training
-(a) Client-batch parallel training.
-(b) Server-hierachical parallel training.
-Figure 5: Multi-client Parallel Training of FL-LLaMA.
-In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation.
-Then, the server splits hidden state $h_{B}$ and sends them back to different clients.
-In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$.
-After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
-In multi-clients settings, the FL-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism and server-hierarchical parallelism.
-Each of them optimizes for different system configurations and resource availability.
-In sequential training, the server interacts with one client at a time in a round-robin manner, which is more suitable for scenarios with limited server capacity.
-In client-batch parallelism, the server waits for data from each client and then concatenate them in a single batch, which is appropriate for single machine with multiple GPUs server scenarios.
-In server-hierarchical parallelism, each client is assigned a separate sub-server for interaction, allowing multiple clients to train in true parallel, which is suitable for multi-machine, multi-GPU server environments.
-3.3.1 Sequential Training
-The client begins by passing hidden states through its input blocks $Blocks_{A}$ and forwards the result to the server.
-The server processes the input through $Blocks_{B}$ and returns the output states.
-The client then computes the loss with $Blocks_{C}$, sends the gradients to the server for model updates, and receives the updated gradients to adjust its input blocks.
-This process continues until all clients have completed their respective training phases for the current round.
-Although this method is inherently slower due to its serialized nature, it significantly reduces the computational burden on the server, making it a suitable choice for scenarios where server-side resources are limited.
-3.3.2 Client-Batch Parallelism
-[…]
-3.3.3 Server-hierarchical Parallelism
-In server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Fig.
-5(b).
-These instances run in parallel, allowing each client to perform local training simultaneously rather than waiting for others to complete.
-Typically after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers.
-This is usually done via weighted averaging of both server-side and client-side model parameters.
-This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
+1-7, the gradient of $W_{n}$ is closely related to the value of $h_{1}$.
+Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations.
+In other words, injecting noise into the forward propagation process can protect the hidden states and their gradients transmitted during the training of the splitting LLMs.
+Inference Accelaration
+The inference time of LLMs is critically important for efficient and fast handling of complex tasks.
+To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
+Attention Mask Compression
+FedSEA-LLaMA transmits only the minimal information necessary for effective model training and inference.
+In particular, it reduces redundancy in attention mask transmissions compared to FL-GLM.
+Since the generation of attention masks remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ results in unnecessary communication overhead.
+Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len) indicating padding locations, thereby significantly decreasing bandwidth usage.
+Collaborative KV Cache Mechanism
+Since the model generates the next token autoregressively based on previously generated tokens, in distributed LLM setups, repeated transmission of the complete context from client to server can drastically increase inference latency as user queries grow longer.
+In centralized LLMs, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers since these states remain unchanged throughout token generation.
+Building upon this insight, FedSEA-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step.
+This approach effectively minimizes data transmission, substantially improving inference efficiency.
+Adaptive Partition
+Erdoğan, Küpçü, and Çiçek (2022) and Pasquini, Ateniese, and Bernaschi (2021) have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations.
+Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
+To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks, thereby achieving an optimal balance between performance and efficiency.
+Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with Blocks_A=1,2,3) and the last blocks (with Blocks_C=1,2,3) of the model.
+The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients.
+Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
+Experiments
+We conduct experiments on NLU, summarization generation, and conversational QA tasks to investigate whether FedSEA-LLaMA can maintain comparable performance to centralized training, while addressing high delays and data privacy concerns through decentralized learning.
+More experiments, including parallel training and training prompts, can be found in the Appendix.
+Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc.
+Acc.
+Acc.
+Acc.
+Acc.
+Acc.
+F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al.
+2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al.
+2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al.
+2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al.
+2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 LLaMA2-7B (Touvron et al.
+2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FL-GLM (Zheng et al.
+2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 - 37.0/11.9 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
+Table 1: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FedSEA-LLaMA and baselines.
+Experimental Settings
+In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FedSEA-LLaMA.
+Dataset
+To evaluate model generalization across diverse NLP tasks, our experiments incorporate three distinct tasks:
+  • The SuperGLUE benchmark (Wang et al.
+2020) is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, from common sense reasoning to fine-grained semantic interpretation.
+  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues (127k+ questions) across seven domains.
+Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
+  • The XSum dataset (Narayan, Cohen, and Lapata 2018) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries.
+It tests the model’s ability to abstract and condense salient information across diverse topics.
+Metrics
+We evaluate model performance using standard task-specific metrics.
+For NLU tasks such as COPA, WSC, etc, we report Accuracy, which measures the proportion of instances where the model’s predicted label exactly matches the ground-truth label.
+For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (Yatskar 2019; You et al.
+2022).
+EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
+For summarization tasks, we use standard ROUGE metrics (Liu et al.
+2021; Chen and Yang 2020; Fang et al.
+2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
+Baselines
+We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (Touvron et al.
+2023), an open-source pre-trained model developed by Meta and FL-GLM.
+We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (Raffel et al.
+2020), $\text{BART}_{\text{Large}}$ (Lewis et al.
+2020), $\text{GLM}_{\text{RoBERTa}}$ (Du et al.
+2022) and ChatGLM-6B (GLM et al.
+2024).
+Parameter Settings
+The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40.
+The implementation of the FedSEA-LLaMA is based on the Flower (Beutel et al.
+2020) federated learning framework, which orchestrates the transfer of intermediate data between clients and the server.
+Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset.
+All other experiments are conducted with FP16 precision and a batch size of 2.
+The random seed is consistently set to 42.
+Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters.
+For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, o_proj.
+Experimental Results
+FedSEA-LLaMA Performance
+In order to validate the effectiveness of FedSEA-LLaMA, we compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 1.
+Experimental results show that FedSEA-LLaMA achieves comparable performance to centralized LLaMA2-7B, demonstrating the feasibility of the framework while significantly reducing GPU memory consumption on the client side.
+The comparison of their GPU memory usage is shown in Figure 4.
+While centralized LLaMA2-7B occupies 28.2G of GPU memory, the FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements.
+Both models were finetuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8.
+Figure 4: Computation load on client side of centralized LLaMA2-7B and FedSEA-LLaMA on ReCoRD dataset.
+(a) Loss Curves for Noise Scales
+(Blocks_A=1, Blocks_C=1)
+(b) Inference Speed Comparison
+(w/ vs. w/o Attention Mask).
+(c) Inference Speed Comparison
+(w/ vs. w/o KV Cache).
+Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention mask. (c)Ablation study on KVcache.
+Effects of Noise Scales
+Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how varying levels of noise affect model performance.
+The results demonstrate that introducing appropriate levels of noise leads to only a slight decrease in the performance of distributed LLMs, while significantly enhancing their security.
+The corresponding experimental results are presented in Figure 5(a) and Table 2.
+Figure 5(a) demonstrates that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ is held constant, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases.
+When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline.
+Table 2 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores.
+Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.
+Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3 0.0 77.48/40.82 77.82/39.24 77.89/41.34 0.02 77.89/38.09 77.87/39.66 78.61/42.39 0.05 74.57/29.17 77.07/38.41 78.95/41.03 0.1 59.11/8.18 72.45/23.71 75.68/34.10
+Table 2: Different noise scales and blocks number of $Blocks_{A}$ analysis(F1a / EM) on MultiRC dataset.
+Model Reverse Attack
+To investigate the impact of split point selection and noise scale on the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5).
+The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5.
+Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2.
+FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57.
+Experiment results of FedSEA-LLaMA is shown in Table  3.
+From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores.
+However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero.
+These findings demonstrate that FedSEA-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
+Noise scale Blocks_A=1 Blocks_A=2 Blocks_A=3 0.0 0.051/0.003 0.056/0.003 0.069/0.016 0.02 0.050/0.002 0.052/0.003 0.084/0.044 0.05 0.050/0.003 0.051/0.002 0.067/0.004
+Table 3: Security analysis (Rouge-2 F1 / BlEU-4) of FedSEA-LLaMA on Record dataset.
+Efficiency Optimization
+Figure 5(b) shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training.
+The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second.
+The context length is 500.
+It can be observed that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
+As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving a speed more than seven times faster.
+Without the KV cache, however, the inference speed decreases significantly as the query length increases, primarily due to the continuously growing number of K and V vectors.
+In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing previously generated K and V vectors.
+This effectively avoids redundant computations.
+When the query length reaches 4000, the inference speed of FedSEA-LLaMA with KV cache is 8.2 times that without it.
+As query length increases further, this advantage continues to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
+Adaptive Splitting
+The choice of model partitioning has a significant impact on the computational burden at the client side.
+To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3).
+This setting reflects realistic constraints where clients possess limited computational resources.
+Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3 1 80.34/44.60 80.63/46.59 80.90/47.11 2 80.41/45.86 80.62/45.96 79.94/45.02 3 81.17/47.22 80.74/46.59 80.20/45.33
+Table 4: Different model partition analysis(F1a/EM) on MultiRC dataset. No Noise is added to the hidden states.
+Table 4 shows F1a/EM scores across different partitioning schemes.
+Performance variations across partitions are minimal, indicating that FedSEA-LLaMA is robust to different split configurations.
+This flexibility allows resource-constrained clients to adjust block allocation based on their downstream task requirements and resource availability, without significant accuracy loss.
+Conclusion
+In this work, we proposed FedSEA-LLaMA, a novel federated split learning framework tailored for large language models.
+By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FedSEA-LLaMA addresses the core limitations of traditional federated split learning approaches.
+Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability.
+This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs.
+In future work, we aim to support more base models and deploy FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
+Appendix A Noise Added to Hidden States
+(a) Loss Curves for Noise Scales
+(Blocks_A=1, Blocks_C=1)
+(b) Loss Curves for Noise Scales
+(Blocks_A=2, Blocks_C=1)
+(c) Loss Curves for Noise Scales
+(Blocks_A=3, Blocks_C=1)
+Figure 6: FedSEA-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
+The magnitude of noise injection significantly influences the stability and convergence speed of model training.
+Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance.
+To quantify this effect, we first fine-tune the FedSEA-LLaMA model on the MultiRC (Khashabi et al.
+2018) with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within the hidden state $h_{A}$.
+According to Figure 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1).
+The average magnitude of the hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$.
+Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance.
+Training is conducted for 5,000 steps with FP16 precision.
+We use LoRA with a rank of 8 and a scaling factor $\alpha=16$.
+The learning rate is 1e-5, and the batch size is 2.
+Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2.
+A horizontal comparison reveals that, under the same noise level, increasing the number of blocks in $Blocks_{A}$ consistently accelerates convergence.
+Table 2 indicates that when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.
+This can be explained by the average magnitude shown in Figure 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR).
+This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
+Figure 7: Avg magnitude of hidden state $h_{A}$ (Blocks_C =1) on MultiRC dataset.
+Appendix B Prompt
+Table 5 illustrates examples of cloze-style questions and corresponding answers derived from each task.
+(1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages.
+(2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise.
+(3)WSC focuses on pronoun disambiguation requiring deep contextual understanding.
+(4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another.
+(5)BoolQ involves yes/no questions answered using evidence from context passages.
+(6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences.
+(7)CB deals with entailment in complex sentences with nuanced reasoning.
+(8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information.
+Table 6 illustrates examples of fine-tuning templates of CoQA and XSum.
 Dataset Task Cloze Question Answers ReCoRD Question answering [passage p] [cloze question q] Answer candidates COPA Causal reasoning “[choice c1]” or“[choice c2]”?
@@
 Table 5: Cloze questions and answers for the 8 SuperGLUE tasks
-3.4 Inference Accelaration
-The inference time of LLMs are critically important for both enterprises and end-users.
-Low-latency inference minimizes user wait times, allowing for efficient and fast handling of complex tasks.
-Distributed architectures partition tensor operations, traditionally executed within GPUs and directly interacting with high-bandwidth memory (HBM), into separate computational units across multiple physical devices.
-Consequently, intermediate data must be transferred over networks, posing substantial latency challenges for the inference of LLMs.
-To enhance the efficiency and usability of distributed LLMs, our proposed FL-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
-3.4.1 Attention Mask Compression
-FL-LLaMA transmits only the minimal information necessary for effective model training and inference.
-One significant improvement arises from reducing redundancy in attention mask transmissions.
-In masked self-attention mechanisms, the attention mask serves two primary purposes: (1) preventing decoders from prematurely accessing subsequent tokens during prediction, and (2) ensuring that padding tokens within input data do not interfere with attention computations involving valid tokens.
-Notably, the generation of attention masks across transformer blocks remains consistent throughout each training iteration and depends solely on dataset padding.
-Consequently, rather than transmitting the entire attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$, which incurs substantial bandwidth latency, the FL-LLaMA framework only sends essential metadata ($seq\_len$, $pad\_len$) indicating the locations of padding tokens versus valid tokens from the client to the server, thus significantly decreasing communication overhead.
-3.4.2 Collaborative KV Cache Mechanism
-Since the model autoregressively generates the next token based on the tokens already generated, the speed of generating new tokens slows down as the input sequence length becomes excessively long (e.g., greater than 1000), due to the increasing number of K and V calculations.
-The application of key-value cache (KV cache) techniques in distributed inference further enhances inference speed.
-Generative LLMs produce outputs autoregressively, meaning that each generated token becomes part of the input sequence used to generate subsequent tokens.
-As user queries grow longer, repeatedly transmitting the complete context—comprising the initial query and previously generated tokens—from client to server can drastically increase inference latency.
-To mitigate this issue, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers.
-Since these states remain unchanged throughout token generation, newly generated tokens reuse cached states without requiring recomputation.
-Building upon this insight, FL-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step.
-This approach effectively minimizes data transmission, substantially improving inference efficiency.
-3.5 Adaptive Partition
-Ege et al.[49] and Dario et al.[50] have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations.
-Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
-To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks in accordance with the demands of downstream tasks and available computational resources, thereby achieving an optimal balance between performance and efficiency.
-Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with Blocks_A=1,2,3) and the last blocks (with Blocks_C=1,2,3) of the model in Sec.
-4.2.6.
-The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients.
-Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
-4 Experiments
-In order to demonstrate the performance of LLaMA within the federated split framework (FL-LLaMA), we conduct experiments using the same benchmarks as those used in LLaMA2-7B model [53].
-Developed by Meta AI, the LLaMA2-7B model serves as a general-purpose language model with strong performance in natural language processing tasks, particularly when fine-tuned for specific domains.
-We employ this architecture to investigate whether FL-LLaMA can maintain comparable task performance to conventional centralized training, while addressing high delays and data privacy concerns through decentralized learning.
-Table 2: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FL-LLaMA and baselines.
-[…]
-4.1 Experimental Settings
-In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FL-LLaMA.
-4.1.1 Dataset
-To evaluate model generalization across diverse NLP tasks, our experiments incorporates three distinct tasks:
-  • 1.
-    The SuperGLUE benchmark [58] is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, which span a broad range of linguistic challenges, from common sense reasoning to fine-grained semantic interpretation.
-Table 1 illustrates examples of cloze-style questions and corresponding answers derived from each task.
-(1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages.
-(2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise.
-(3)WSC focuses on pronoun disambiguation requiring deep contextual understanding.
-(4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another.
-(5)BoolQ involves yes/no questions answered using evidence from context passages.
-(6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences.
-(7)CB deals with entailment in complex sentences with nuanced reasoning.
-(8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information.
-To unify these tasks under a single modeling framework, we adopt the formulation used in GLM [56], representing each instance as a blank infilling problem.
-Given a labeled pair $(x,y)$, the input $x$ is converted into a masked cloze question $q(x)$ using a placeholder token $[M]$, while the output $y$ serves as the target answer $a(y)$.
-  • 2.
-    The CoQA dataset [59] is designed for the task of conversational question answering(QA), focusing on evaluating contextual coherence in multi-turn dialogues.
-It includes over 8,000 dialogues and more than 127,000 questions drawn from seven diverse domains.
-Answers are provided in free-form text, with nearly half of the questions requiring coreference resolution (e.g., pronoun disambiguation) and pragmatic reasoning.
-This setup closely mirrors real-world conversational settings, aiming to assess a model’s ability to understand text passages and respond to a series of interrelated questions.
-  • 3.
-    The XSum dataset [60] is designed for the task of abstractive summarization, specifically evaluating a model’s ability to perform extreme compression of salient information from news articles.
-It consists of BBC news articles paired with manually written summaries, covering a broad range of topics.
-A key challenge of XSum lies in its requirement for the model to generate a single-sentence summary for each article, placing significant demands on the model’s ability to abstract and condense information effectively.
-4.1.2 Metrics
-Since NLU tasks are reformulated as blank infilling problems, model performance is assessed based on the probability assigned to the ground-truth answer $a(y)$.
-For datasets such as RTE, BoolQ, WiC, CB, and MultiRC where the correct answer is often a single token, the evaluation score is derived from the logit corresponding to that token.
-Specifically, we define the probability as follows:
-$p(y|x)=\frac{p(a(y)|q(x))}{\sum_{y^{{}^{\prime}}\in Y}p(a(y^{{}^{\prime}})|q(x))},$
-where $Y$ represents the set of possible ground-truth labels.
-In contrast, datasets like ReCoRD, COPA, and WSC require multi-token answers.
-In these cases, we evaluate model performance by summing the log-probabilities of each token in the predicted answer, computed as:
-$s(y|x)=\sum_{t=1}^{|L_{y}|}\text{log}\text{P}(y_{t}|y_{1}...y_{t-1},x;\theta).$
-For conversational QA tasks, we utilize two metrics: Exact Match (EM) and F1 score to compare the model performance, which evaluate the accuracy of predicted answers against the ground truth [61, 62].
-Exact Match (EM) measures the percentage of predictions that match the ground truth answers exactly, accounting for variations in punctuation and case.
-The F1 score is the harmonic mean of precision and recall, providing a more nuanced view of performance by rewarding partial overlaps between the predicted and reference answers.
-For summarization tasks, we adopt standard ROUGE metrics [63, 64, 65], i.e., ROUGE-1 and ROUGE-2, which are widely used to measure content overlap between generated summaries and reference texts.
-ROUGE-1 evaluates the unigram overlap, while ROUGE-2 measures the bigram overlap, providing insights into both content coverage and fluency.
-4.1.3 Baselines
+Dataset Task Fine-tuning template
+CoQA conversational QA
+<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
+You need to answer my question base on the provided story.<</SYS>><</SYS>>
+The story is {story}.My question is {question}.[/INST]{answer}</s></s>
+XSum summarization generation
+<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
+You need to summarize an abstract base on the provided passage.<</SYS>><</SYS>>
+The passage is {passage}.[/INST]{summarization} </s></s>
+Table 6: Fine-tuning templates for CoQA and XSum
+Appendix C Parallel Training
+Strategy Sequential training    Client-batch training    Server-hierachical training      Centralized
+num. of clients 2 2 4 8 2 4 8 -
 […]
 Table 7: Comparison of training time between different training strategies on ReCoRD dataset.
-Strategy Sequential training Client-batch training Server-hierachical training   Centralized
-num. of clients 2 2 4 8 2 4 8 -
-time(s) 689.2$\pm$5.7 630.0$\pm$1.8 542.2$\pm$3.5 488.6$\pm$0.4 348.2$\pm$3.8 173.53$\pm$3.0 86.4$\pm$0.6 381.9$\pm$2.4
-4.1.4 Parameter Settings
-The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40.
-The implementation of the FL-LLaMA is based on the Flower [67] federated learning framework, which orchestrates the transfer of intermediate data between clients and the server.
-Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset.
-All other experiments are conducted with FP16 precision and a batch size of 2.
-The random seed is consistently set to 42.
-Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters.
-4.2 Experimental Results
-4.2.1 FL-LLaMA Performance
-In order to validate the effectiveness of FL-LLaMA, we compared the performance of FL-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 2.
-Experimental results show that FL-LLaMA and centralized LLaMA2-7B exhibit no significant performance differences, demonstrating the feasibility of the framework.
-Not only does FL-LLaMA perform comparably to centralized LLaMA2-7B, but the GPU memory consumption on the client side is also significantly reduced compared to centralized LLaMA2-7B.
-The comparison of their GPU memory usage is shown in Fig.
-6.
-While centralized LLaMA2-7B occupies 27.9G of GPU memory, the FL-LLaMA client requires only 6G, representing a reduction of up to 82% in memory requirements.
-This reduction enables enterprises or departments that possess large amounts of high-quality private datasets but have limited computational resources to participate in federated training of LLMs.
-Figure 6: Computation load on client side of centralized LLaMA2-7B and FL-LLaMA on ReCoRD dataset.
-4.2.2 Parallel Training
-For communication cost analysis, we randomly selected 1,000 samples from the ReCoRD dataset.
-In our experimental setup, all training used a batch size of 1 and adopted FP16 training to improve computational efficiency.
-For LoRA, the decomposition rank $r$ is set to 8, and the scaling factor $\alpha$ is configured to 16, focusing specifically on the attention modules during fine-tuning.
-Additionally, the model is partitioned with Blocks_A=1 and Blocks_C=1.
-Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time, as summarized in Tab.
-3.
-The results show that sequential training takes the longest time, followed by client-batch training, and server hierarchical takes the shortest time.
-However, when the server can accommodate a larger batch size, increasing the number of clients in client-batch training effectively increases the batch size and thus reduces overall training time.
-Figure 7: Avg magnitude of hidden state $h_{A}$ (Blocks_C =1) on MultiRC dataset.
-(a) Loss Curves for Noise Scales
-(Blocks_A=1, Blocks_C=1)
-(b) Loss Curves for Noise Scales
-(Blocks_A=2, Blocks_C=1)
-(c) Loss Curves for Noise Scales
-(Blocks_A=3, Blocks_C=1)
-Figure 8: FL-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
-4.2.3 Effects of Noise Scales
-Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how varying levels of noise affect model performance.
-The results demonstrate that introducing appropriate levels of noise leads to only a slight decrease in the performance of distributed LLMs, while significantly enhancing their security.
-The magnitude of noise injection significantly influences the stability and convergence speed of model training.
-Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance.
-To quantify this effect, we first fine-tuned the FL-LLaMA model on the MultiRC [68] with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within hidden state $h_{A}$.
-According to Fig.
-7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1).
-The average magnitude of hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$.
-Based on this observation, we subsequently injected noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance.
-Training is conducted for 5,000 steps with FP16 precision.
-We used LoRA with a rank of 8 and a scaling factor $\alpha=16$.
-The learning rate is 1e-5, and the batch size is 2.
-The corresponding experimental results are presented in Fig.
-8(a)-8(c) and Tab.
-4.
-Fig.
-8(a)-8(c) demonstrate that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ are held constant, the convergence speed of FL-LLaMA decreases as the magnitude of injected noise increases.
-When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline.
-A horizontal comparison reveals that, under the same noise level, increasing the number of blcoks in $Blocks_{A}$ consistently accelerates convergence.
-Tab.
-4 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores.
-Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.
-This can be explained by the average magnitude shown in Fig.
-7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR).
-This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
-Table 4: Different noise scales and blocks number of $Blocks_{A}$ analysis(F1a / EM) on MultiRC dataset.
-Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3
-0.0 77.48/40.82 77.82/39.24 77.89/41.34
-0.02 77.89/38.09 77.87/39.66 78.61/42.39
-0.05 74.57/29.17 77.07/38.41 78.95/41.03
-0.1 59.11/8.18 72.45/23.71 75.68/34.10
-4.2.4 Model Reverse Attack
-To investigate the impact of split point selection and noise scale on the security of FL-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5).
-The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5.
-Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2.
-FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57.
-Experiment results of FL-LLaMA is shown in Tab.
-5.
-From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores.
-However, under the FL-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero.
-Tab.
-6 reports the downstream performance of $client_{0}$ after multi-client training, showing that secret attack allows training to proceed as usual without disrupting the federated process.
-These findings demonstrate that FL-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
-Table 5: Security analysis (Rouge-2 F1 / BlEU-4) of FL-LLaMA on Record dataset.
-Noise scale Blocks_A=1 Blocks_A=2 Blocks_A=3
-0.0 0.051/0.003 0.056/0.003 0.069/0.016
-0.02 0.050/0.002 0.052/0.003 0.084/0.044
-0.05 0.050/0.003 0.051/0.002 0.067/0.004
-Table 6: Performance (EM) under different noise scales and blocks number of $Blocks_{A}$ on Record dataset.
-Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3
-0.0 77.4 79.8 78.9
-0.02 79.1 79.9 79.6
-0.05 79.6 79.4 79.2
-4.2.5 Efficiency Optimization
-Figure 9: FL-LLaMA: Inference Speed Comparison (w/ vs. w/o Attention Mask).
-Fig.
-9 shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training.
-The bar chart represents the average time spent per token in FL-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second by FL-LLaMA.
-It can be observed from the figure that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
-Figure 10: FL-LLaMA: Inference Speed Comparison(w/ vs. w/o KV Cache).
-Fig.
-10 illustrates the effect of the KV cache technique on improving inference speed for different query lengths in distributed inference.
-The experimental results indicate that without the KV cache, the inference speed of FL-LLaMA significantly decreases as the query length increases, primarily due to the continuously increasing number of K and V vectors computed at each step in the self-attention mechanism.
-In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing the previously generated K and V vectors, thereby effectively avoiding redundant computations.
-When the query length reaches 4000, the inference speed of FL-LLaMA with KV cache is 8.2 times that without it, and as the query length increases further, this advantage will continue to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
-4.2.6 Adaptive Splitting
-The choice of model partitioning has a significant impact on the computational burden at the client side.
-To evaluate whether FL-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3).
-This setting reflects realistic constraints where clients possess limited computational resources.
-All models are trained for 8,000 steps using FP16 precision.
-LoRA is applied with a rank of 16 and a scaling factor of $\alpha=32$.
-The training used a batch size of 2 and a learning rate of 2e-5.
-Table 7: Different model partition analysis(F1a/EM) on MultiRC dataset. No Noise is added to the hidden states.
-Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3
-1 80.34/44.60 80.63/46.59 80.90/47.11
-2 80.41/45.86 80.62/45.96 79.94/45.02
-3 81.17/47.22 80.74/46.59 80.20/45.33
-Tab.
-7 presents F1a/EM scores under various partitioning schemes.
-The highest F1a/EM scores (81.17/47.22) emerge when Blocks_A=3 and Blocks_C = 1, indicating that allocating more early blocks to the client side can slightly enhance task performance.
-Overall, the performance differences across partitions remain small, suggesting that FL-LLaMA is relatively insensitive to how the model is split.
-From a practical standpoint, these findings confirm that clients with constrained resources can flexibly adjust how many blocks they store without substantially degrading accuracy.
-In scenarios where client-side memory or compute capacities are limited, reducing Blocks_A and Blocks_C remains viable, given that FL-LLaMA maintains robust performance across a range of partitions.
-Building upon partitioning experiment, we introduced controlled noise into the hidden states produced by $Blocks_{A}$.
-Specifically, Gaussian noise with a scale of 0.02 is added during forward propagation before sending the hidden states to the server.
-Tab.
-8 shows that while there is a slight drop compared to the noise-free scenario, the overall performance remains robust, indicating that moderate perturbations in the hidden states do not severely impact convergence or accuracy.
-Futher, the experimental results show that the higher the number of $Blocks_{A}$, the smaller the performance degradation of FL-LLaMA on downstream tasks.
-This indicates that the resistance to noise is higher, which verifies that the signal-to-noise ratio has a significant impact on the performance of the model.
-Table 8: Different model partitions analysis(F1a/EM) on MultiRC dataset. Noise (scale=0.02) is added to the states $h_{A}$.
-Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3
-1 79.44/38.41 79.58/38.61 79.45/39.66
-2 80.10/41.55 80.14/38.61 79.99/40.40
-3 79.90/40.29 80.29/43.02 80.44/42.18
-5 Conclusion
-In this work, we proposed FL-LLaMA, a novel federated split learning framework tailored for large language models.
-By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FL-LLaMA addresses the core limitations of traditional federated split learning approaches.
-Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FL-LLaMA demonstrates significant advances in security, efficiency, and adaptability.
-This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs.
-In future work, we aim to support more basement models and deploy FL-LLaMA framework in scenarios where data privacy is a critical concern.
+(a) Client-batch parallel training.
+(b) Server-hierachical parallel training.
+Figure 8: Multi-client Parallel Training of FedSEA-LLaMA.
+In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation.
+Then, the server splits hidden state $h_{B}$ and sends them back to different clients.
+In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$.
+After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
+In multi-client settings, the FedSEA-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism, and server-hierarchical parallelism, each designed for different system configurations and resource availabilities.
+Sequential Training
+The client begins by passing hidden states through its input blocks $Blocks_{A}$ and forwards the result to the server.
+The server processes the input through $Blocks_{B}$ and returns the output states.
+The client then computes the loss with $Blocks_{C}$, sends the gradients to the server for model updates, and receives the updated gradients to adjust its input blocks.
+This process continues until all clients have completed their respective training phases for the current round.
+Although this method is inherently slower due to its serialized nature, it significantly reduces the computational burden on the server, making it a suitable choice for scenarios where server-side resources are limited.
+Client-batch Parallelism
+The client-batch parallel approach enhances training throughput by expanding the effective batch size at the server, as shown in Figure  8(a).
+During each training round, the server collects intermediate hidden states from all participating clients.
+If each of the $M$ clients processes a local batch of size $b$, the hidden states sent to the server will each have the shape $(b,\textit{seq\_length},\textit{hidden\_size})$.
+The server concatenates these hidden states along the batch dimension, forming a combined tensor of shape $(b\times M,\textit{seq\_length},\textit{hidden\_size})$.
+This aggregated batch is then used to perform a joint forward and backward pass on the server model, allowing it to compute gradients more efficiently and update shared parameters.
+Server-hierarchical Parallelism
+In a server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Figure  8(b).
+These instances run in parallel, enabling simultaneous local training for all clients.
+Typically, after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers.
+This is usually done via weighted averaging of both server-side and client-side model parameters.
+This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
+For communication cost analysis, we randomly select 1,000 samples from the ReCoRD dataset.
+All training uses a batch size of 1 with FP16 precision for efficiency.
+LoRA is applied with a rank of 8 and a scaling factor of 16, targeting attention modules during fine-tuning.
+The model is partitioned with Blocks_A=1 and Blocks_C=1.
+Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time in Table 7.
+Results show that sequential training is the slowest, client-batch training is faster, and server hierarchical training is the fastest.
+Notably, when the server can handle larger batches, increasing the number of clients in client-batch training effectively reduces training time.
+Appendix D Parameter Setting
+FedSEA-LLaMA and Baselines
+Different hyperparameters are explored to obtain the best performance of FedSEA-LLaMA and Centralized LLaMA.
+Specifically, we tune the LoRA rank (r) and scaling factor (alpha) as pairs, trying combinations including (4, 32), (8, 32), (16, 32), (32, 32), and (8, 16).
+For the learning rate (lr), values of 1e-5, 2e-5, and 3e-5 are tested.
+The final parameter settings were selected based on the model’s validation performance.
+The final parameters are shown in Table 8.
+Table 8: Hyperparameters for FedSEA-LLaMA and Baselines
+CoQA XSum MultiRC Record
+FedSEA-LLaMA Lora r & alpha $(4,32)$ $(16,32)$ $(32,32)$ $(32,32)$
+lr 1e-4 3e-5 1e-5 2e-5
+Centralized-LLaMA Lora r & alpha $(8,16)$ $(32,32)$ $(32,32)$ $(8,32)$
+lr 5e-5 3e-5 1e-5 3e-5
+Different Noise Scales and Layers of $Block_{A}$ Experiment
+We use a LoRA rank of 8 with a scaling factor of 16, a learning rate of 1e-5, and FP16 precision to investigate the effects of different noise scales and the number of blocks in $Blocks_{A}$ on downstream performance.
+Ablation Study on Attention Mask
+The context length is set to 500, and the precision is set to FP16.
+Ablation Study on KV Cache
+The precision is set to FP16. Different context lengths are set to investigate the inference speed of FedSEA-LLaMA and FL-GLM.
+Different Model Partition
+All experimental results were obtained after 8,000 training steps with LoRA rank set to 8, scaling factor set to 16, a batch size of 1, FP16 precision, and a learning rate of 1e-5.

### v3

@@
 Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements.
-Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups of training and inference.
+Experiments on natural language understanding, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups in training and inference.
 Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FedSEA-LLaMA in security and adaptability.
+Codehttps://github.com/TAP-LLM/SplitFedLLM/tree/main/FedSEA-LLaMA
+Extended versionhttps://arxiv.org/abs/2505.15683
 Introduction
@@
 […]
-The distributed nature of private data across various silos and the significant computational demands of LLMs present substantial challenges for their deployment in federated environments.
-Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprises servers (Huang et al.
-2024) or medical institutions (Qi et al.
-2024; Yuan et al.
+The distributed private data silos and the significant computational demands of LLMs pose substantial challenges for their deployment in federated environments.
+Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprise servers (Huang et al.
+2024), or medical institutions (Yuan et al.
 2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al.
 2024).
-Federated Learning (FL) (Stremmel and Singh 2021; Li et al.
-2020) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation.
+Federated Learning (FL) (Stremmel and Singh 2021) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation.
 They often leverage Adapters (Houlsby et al.
 2019) or LoRA (Hu et al.
-2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to achieve fine-tuning LLMs.
+2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to fine-tune LLMs.
 However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible.
-Therefore, how to effectively deploy LLMs within FL frameworks while ensuring privacy and scalability remains an open research problem.
+Therefore, how to effectively deploy LLMs in FL frameworks while ensuring privacy and scalability in resource-constrained scenarios remains an open research problem.
 To address this, transformer-based federated splitting models (Zheng et al.
 2024; Su et al.
-2024) have been proposed to adapt LLMs fine-tuning to the resource-constrained scenarios.
+2024) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios.
 Titanic (Su et al.
@@
 FL-GLM (Zheng et al.
-2024) offloads most model parameters to resource-ample server while retaining the input and output layer on client to ensure data privacy.
-As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on client, while the middle blocks on the server.
+2024) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client.
+As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on the client, while the middle blocks on the server.
 The forward propagation (FP) path is client-to-server-to-client with peer-to-peer key encryption.
-After client calculates loss, gradients backpropagate(BP) from client-to-server-to-client with the same encryption.
-Although FL-GLM provides a potentially viable solution for clients resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al.
+After the client calculates loss, gradients backpropagate (BP) from client-to-server-to-client with the same encryption.
+Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al.
 2023).
@@
 In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated split framework based on LLaMA2, as shown in Figure 1b.
-To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks, and send these noised states to the server for further processing.
+To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks and send these noised states to the server for further processing.
 Through both gradient analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages.
 To Reduce Communication Cost, we introduce the attention-mask compression and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of auto-regressive inference.
-By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamically generated local data, reducing communication from megabytes(MB) to bytes(B).
+By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamic attention-mask generation, reducing communication from megabytes (MB) to bytes (B).
 To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources.
 This adaptability enables an optimal balance between performance and efficiency.
-Experiments on NLU, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2, and achieves up to 8× inference speedups.
+Experiments on natural language understanding (NLU), summarization, and conversational QA show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× inference speedups.
 Through inference acceleration, the average token generation time of FedSEA-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the growing length of context.
 Additional analysis of privacy attacks and various partition points further highlights FedSEA-LLaMA’s effectiveness in terms of security and adaptability.
-[^1] Our code is available after the submission is accepted
 The innovations in this paper are as follows:
   • We design a federated split framework tailored for resource-limited scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
-  • We leverage collaborative KV caching and dynamic attention mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
-  • Experimental results on different three tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency and adaptability of our framework.
+  • We leverage collaborative KV caching and dynamic attention-mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
+  • Experimental results on three different tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency, and adaptability of our framework.
 Related Works
@@
 2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices.
-Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points and susceptibility to privacy threats like embedding gradient attacks.
+Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points, and susceptibility to privacy threats like embedding gradient attacks.
 Recently, Su et al.
@@
 Zheng et al.
-(2024) propose FL-GLM to place the input and output blocks locally on client devices, while remaining primary model parameters are centralized on a server with ample computational resources.
-And then, they employ key-encryption during client-server communication to prevent reverse engineering attacks from peer clients.
+(2024) propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources.
+And then, they employ key encryption during client-server communication to prevent reverse engineering attacks from peer clients.
 But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
-Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
+Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points, and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
 Model
-In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration and adaptive partitioning.
-Followed FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks and the last q{q} blocks, while the remaining k{k} blocks, comprising the majority of the model parameters, are hosted on the server.
-Next, we design a Gaussian noise injection mechanism applied to the forward-pass hidden states to ensure data privacy during transmission between the client and server.
-Third, we accelerate inference by leveraging the attention mask compression and collaborative KV cache mechanisms.
+In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration, and adaptive partitioning.
+Following FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks($Blocks_{A}$) and the last q{q} blocks($Blocks_{C}$), while the remaining k{k} blocks($Blocks_{B}$), comprising the majority of the model parameters, are hosted on the server.
 Privacy Preservation
@@
 […]
-Threat Model Defination
-We design an extreme attack scenario, inspired by Pasquini, Ateniese, and Bernaschi (2021), to evaluate the robustness of FedSEA-LLaMA.
-Specifically, we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server.
+Threat Model Definition
+We design an extreme attack scenario to evaluate the robustness of FedSEA-LLaMA.
+Inspired by Pasquini, Ateniese, and Bernaschi (2021), we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server.
 The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally.
 The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client.
-During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without their awareness.
+During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without awareness.
 Concretely, we consider an adversary that has full knowledge of the client-side model architecture and the number of layers deployed on the client, but does not have access to the actual model parameters.
@@
 Afterward, $h_{B}$ is split and sent back to the respective clients, allowing training to proceed as usual without disrupting the federated process.
-Defense with Gaussion Noise Injection
+Defense with Gaussian Noise Injection
 We explore two defense strategies to protect the transmission in FedSEA-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states.
@@
 Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation.
+Without loss of generality, we apply noise perturbation to $h_{1}$ in the theoretical proof, although in practice we apply noise perturbation to $h_{A}$.
+The proof will be similar in both cases.
 Figure 3b illustrates the schematic of noise injection during the forward propagation of LLMs.
-Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer and $W_{n+1}$ for the weight of output linear layer.
+Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer, and $W_{n+1}$ denote the weight of the output linear layer.
 In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
@@
 According to Eq.
-1-7, the gradient of $W_{n}$ is closely related to the value of $h_{1}$.
-Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations.
+1-11, the gradient of $W_{n}$ is closely related to the value of $h_{1}$.
+Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into the gradients of parameters, thereby influencing subsequent gradient computations.
 In other words, injecting noise into the forward propagation process can protect the hidden states and their gradients transmitted during the training of the splitting LLMs.
-Inference Accelaration
-The inference time of LLMs is critically important for efficient and fast handling of complex tasks.
-To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
-Attention Mask Compression
+Inference Acceleration
+The inference time of LLMs is crucial for the efficient and timely execution of complex tasks.
+To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention-mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
+Attention-Mask Compression
 FedSEA-LLaMA transmits only the minimal information necessary for effective model training and inference.
-In particular, it reduces redundancy in attention mask transmissions compared to FL-GLM.
-Since the generation of attention masks remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ results in unnecessary communication overhead.
-Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len) indicating padding locations, thereby significantly decreasing bandwidth usage.
+In particular, it reduces redundancy in attention-mask transmissions compared to FL-GLM.
+Since the generation of attention-mask remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention-mask tensors $A_{mask}\in[0,-\infty]^{batch\_size\times seq\_len\times seq\_len}$ results in unnecessary communication overhead during training and inference.
+Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len), indicating padding locations, thereby significantly decreasing bandwidth usage.
 Collaborative KV Cache Mechanism
-Since the model generates the next token autoregressively based on previously generated tokens, in distributed LLM setups, repeated transmission of the complete context from client to server can drastically increase inference latency as user queries grow longer.
-In centralized LLMs, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers since these states remain unchanged throughout token generation.
-Building upon this insight, FedSEA-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step.
-This approach effectively minimizes data transmission, substantially improving inference efficiency.
+In traditional distributed inference of LLMs, each inference step involves transmitting all previously generated tokens from the client to the server.
+This design leads to two major inefficiencies: (1) the computational overhead increases linearly with the growing input length, since the attention mechanism must recompute key and value (KV) representations for all past tokens; and (2) the repeated transmission of the entire context between client and server introduces substantial communication latency, especially for long sequences.
+To address these issues, we propose a collaborative KV cache mechanism tailored for distributed inference.
+Instead of recomputing and retransmitting all past tokens at each step, our approach caches the attention KV states locally on both the client and server sides.
+Since these KV states remain static across the autoregressive generation process, subsequent tokens can reuse the cached representations.
+Built on this mechanism, FedSEA-LLaMA enables highly efficient distributed inference by significantly reducing both computational and communication overhead.
+Experimental results show that this strategy leads to substantial improvements in inference speed in scenarios with long user queries.
 Adaptive Partition
-Erdoğan, Küpçü, and Çiçek (2022) and Pasquini, Ateniese, and Bernaschi (2021) have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations.
-Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
 To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks, thereby achieving an optimal balance between performance and efficiency.
@@
 The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients.
-Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
+Furthermore, it corroborates the efficiency of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
 Experiments
-We conduct experiments on NLU, summarization generation, and conversational QA tasks to investigate whether FedSEA-LLaMA can maintain comparable performance to centralized training, while addressing high delays and data privacy concerns through decentralized learning.
+We evaluate FedSEA-LLaMA on NLU, summarization, and conversational QA to assess its performance against centralized training while tackling latency and privacy via decentralized learning.
 More experiments, including parallel training and training prompts, can be found in the Appendix.
@@
 2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al.
-2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 LLaMA2-7B (Touvron et al.
-2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FL-GLM (Zheng et al.
-2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 - 37.0/11.9 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
+2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (Zheng et al.
+2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (Touvron et al.
+2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
 Table 1: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FedSEA-LLaMA and baselines.
 Experimental Settings
-In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FedSEA-LLaMA.
+In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines, and parameter settings for FedSEA-LLaMA.
 Dataset
 To evaluate model generalization across diverse NLP tasks, our experiments incorporate three distinct tasks:
-  • The SuperGLUE benchmark (Wang et al.
-2020) is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, from common sense reasoning to fine-grained semantic interpretation.
-  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues (127k+ questions) across seven domains.
+  • The SuperGLUE benchmark (Wang et al. 2020) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
+  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains.
 Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
@@
 EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
-For summarization tasks, we use standard ROUGE metrics (Liu et al.
-2021; Chen and Yang 2020; Fang et al.
+For summarization tasks, we use standard ROUGE metrics (Lin 2004; Liu et al.
+2021; Fang et al.
 2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
@@
 The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40.
-The implementation of the FedSEA-LLaMA is based on the Flower (Beutel et al.
-2020) federated learning framework, which orchestrates the transfer of intermediate data between clients and the server.
+The implementation is based on the Flower (Beutel et al.
+2020) framework, which orchestrates the transfer of intermediate data between clients and the server.
 Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset.
@@
 Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters.
-For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, o_proj.
+For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, and o_proj.
 Experimental Results
 FedSEA-LLaMA Performance
-In order to validate the effectiveness of FedSEA-LLaMA, we compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 1.
-Experimental results show that FedSEA-LLaMA achieves comparable performance to centralized LLaMA2-7B, demonstrating the feasibility of the framework while significantly reducing GPU memory consumption on the client side.
+We compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks, including NLU, summarization generation, and conversational QA, as shown in Table 1.
+Experimental results show that FedSEA-LLaMA achieves comparable performance to LLaMA2-7B, with metrics approaching or matching across multiple tasks, demonstrating its effectiveness.
+Meanwhile, compared to other baselines, FedSEA-LLaMA achieves better performance on datasets such as Record, MultiRC, and Xsum, but performs poorly on the RTE dataset.
+This is speculated to be related to the distribution difference between RTE and the data used for LLaMA2 pre-training, as well as suboptimal hyperparameters.
 The comparison of their GPU memory usage is shown in Figure 4.
-While centralized LLaMA2-7B occupies 28.2G of GPU memory, the FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements.
-Both models were finetuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8.
+The result shows that FedSEA-LLaMA significantly reduces GPU memory consumption on the client side.
+Both models were fine-tuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8.
+While centralized LLaMA2-7B occupies 28.2G of GPU memory, FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements.
 Figure 4: Computation load on client side of centralized LLaMA2-7B and FedSEA-LLaMA on ReCoRD dataset.
@@
 (b) Inference Speed Comparison
-(w/ vs. w/o Attention Mask).
+(w/ vs. w/o Attention-Mask).
 (c) Inference Speed Comparison
 (w/ vs. w/o KV Cache).
-Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention mask. (c)Ablation study on KVcache.
+Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention-mask. (c)Ablation study on KV cache.
 Effects of Noise Scales
-Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how varying levels of noise affect model performance.
-The results demonstrate that introducing appropriate levels of noise leads to only a slight decrease in the performance of distributed LLMs, while significantly enhancing their security.
+Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how noise levels affect model performance.
 The corresponding experimental results are presented in Figure 5(a) and Table 2.
-Figure 5(a) demonstrates that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ is held constant, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases.
+Figure 5(a) demonstrates that when the numbers of blocks in $Blocks_{A}$ and $Blocks_{C}$ are consistent, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases.
 When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline.
@@
 Model Reverse Attack
-To investigate the impact of split point selection and noise scale on the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5).
-The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5.
-Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2.
+To assess how split points and noise scale affect the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth ($Blocks_{A}$ = 1, 2, 3) and noise (scale = 0, 0.02, 0.5).
+Attacks use full-parameter fine-tuning (lr=2e-5), with reconstruction quality measured by BLEU-4 and ROUGE-2.
 FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57.
-Experiment results of FedSEA-LLaMA is shown in Table  3.
+Experiment results of FedSEA-LLaMA are shown in Table  3.
 From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores.
-However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero.
+However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLM blocks in addition to the embedding layer, all attack metrics approach zero.
 These findings demonstrate that FedSEA-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
@@
 Efficiency Optimization
-Figure 5(b) shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training.
-The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second.
-The context length is 500.
-It can be observed that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
-As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving a speed more than seven times faster.
-Without the KV cache, however, the inference speed decreases significantly as the query length increases, primarily due to the continuously growing number of K and V vectors.
-In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing previously generated K and V vectors.
+Figure 5(b) shows a comparison of the time consumption between transferring the attention-mask and transferring only the actual input length during distributed training.
+The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second(context length=500).
+We can see that transferring the compressed information of the attention-mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
+As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving over 7× speedup.
+Without the KV cache, the inference speed decreases significantly with longer queries, due to the increasing number of K and V vectors.
+In contrast, with the KV cache technique, the client only needs to send the intermediate vector of the current single token to the server, while reusing previously generated K and V vectors.
 This effectively avoids redundant computations.
@@
 The choice of model partitioning has a significant impact on the computational burden at the client side.
-To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3).
+To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different numbers of blocks in $Blocks_{A}$ and $Blocks_{C}$ ($Blocks_{A}$ = 1, 2, 3 and $Blocks_{C}$ = 1, 2, 3).
 This setting reflects realistic constraints where clients possess limited computational resources.
-Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3 1 80.34/44.60 80.63/46.59 80.90/47.11 2 80.41/45.86 80.62/45.96 79.94/45.02 3 81.17/47.22 80.74/46.59 80.20/45.33
-Table 4: Different model partition analysis(F1a/EM) on MultiRC dataset. No Noise is added to the hidden states.
 Table 4 shows F1a/EM scores across different partitioning schemes.
@@
 This flexibility allows resource-constrained clients to adjust block allocation based on their downstream task requirements and resource availability, without significant accuracy loss.
+$Blocks_{A}$ $Blocks_{C}$ = 1 $Blocks_{C}$ = 2 $Blocks_{C}$ = 3 1 80.34/44.60 80.63/46.59 80.90/47.11 2 80.41/45.86 80.62/45.96 79.94/45.02 3 81.17/47.22 80.74/46.59 80.20/45.33
+Table 4: Different model partition analysis(F1a/EM) on MultiRC dataset. No Noise is added to the hidden states.
 Conclusion
-In this work, we proposed FedSEA-LLaMA, a novel federated split learning framework tailored for large language models.
+In this work, we propose FedSEA-LLaMA, a novel federated split learning framework tailored for LLMs.
 By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FedSEA-LLaMA addresses the core limitations of traditional federated split learning approaches.
-Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability.
+Through secure end-to-end propagation with noise injection, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability.
 This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs.
-In future work, we aim to support more base models and deploy FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
+In future work, we aim to support more base models and deploy the FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
+Acknowledgements
+This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No.
+62406013,No.62388101, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-24-034), and the Fundamental Research Funds for the Central Universities.
 Appendix A Noise Added to Hidden States
@@
 Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2.
+To assess the sensitivity to the noise scale when $Blocks_{A}$=3, we evaluate performance across a range of noise scales (from 0.1 to 0.4).
+As shown in Table 5, model performance decreases as the noise scale increases.
+This suggests that while moderate noise injection is tolerable, excessive noise harms representation quality.
+Noise Scale F1a EM
+0.1 75.68 34.10
+0.15 71.21 26.82
+0.2 59.08 16.47
+0.3 55.79 7.45
+0.4 57.68 6.09
+Table 5: Performance under different noise scales when $Blocks_{A}$=3
 A horizontal comparison reveals that, under the same noise level, increasing the number of blocks in $Blocks_{A}$ consistently accelerates convergence.
@@
 Appendix C Parallel Training
-Strategy Sequential training    Client-batch training    Server-hierachical training      Centralized
-num. of clients 2 2 4 8 2 4 8 -
 […]
@@
 Notably, when the server can handle larger batches, increasing the number of clients in client-batch training effectively reduces training time.
-Appendix D Parameter Setting
+Appendix D Efficiency Comparison between FedSEA-LLaMA and FL-GLM
+For the measurements of inference speed, inference memory usage, and per-token communication volume, we used the following prompt template:
+{B_INST}{B_SYS}System: You are a helpful assistant. Always answer as helpfully as possible.{E_SYS}User: what can you do?{E_INST}
+The total prompt length was 44 tokens.
+All inference experiments were conducted in FP16 precision with a maximum output length of 128 tokens.
+During inference, the client held one layer of model A and one layer of model C.
+For the training memory usage experiments, we used the MultiRC dataset.
+Both model A and model C contained one layer.
+Unless otherwise specified, the default configuration was as follows: learning rate $1\times 10^{-5}$, LoRA rank =32=32, LoRA scaling factor =32=32, random seed =64=64, FP16 precision, and batch size =2=2.
+It is important to note that our training implementation actively clears unused variables in real time to minimize GPU memory usage.
+Therefore, the reported memory values correspond to the maximum observed usage within a one-minute monitoring window.
+All experiments were repeated three times with different random seeds (42, 64, and 128) to ensure the robustness and reproducibility of the results.
+The comparison results are shown in the Table 9.
+Framework
+Training Memory Usage
+Inference Memory Usage
+per-token communication
+volume (Bytes)
+Inference tokens
+per second
+FL-GLM 9810±0.00 6886±0.00 context length N × 10,335 0.75±0.03
+FedSEA-LLaMA 5078±0.00 3674.00±0.00 8494 0.34±0.01
+Table 9: Efficiency Comparison between FedSEA-LLaMA and FL-GLM
+Appendix E Parameter Setting
 FedSEA-LLaMA and Baselines
@@
 The final parameters are shown in Table 10.
-Table 8: Hyperparameters for FedSEA-LLaMA and Baselines
 CoQA XSum MultiRC Record
@@
 lr 5e-5 3e-5 1e-5 3e-5
+Table 10: Hyperparameters for FedSEA-LLaMA and Baselines
 Different Noise Scales and Layers of $Block_{A}$ Experiment
 We use a LoRA rank of 8 with a scaling factor of 16, a learning rate of 1e-5, and FP16 precision to investigate the effects of different noise scales and the number of blocks in $Blocks_{A}$ on downstream performance.
-Ablation Study on Attention Mask
+Ablation Study on Attention-Mask
 The context length is set to 500, and the precision is set to FP16.

### v4

@@
 Introduction
-While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (Kaplan et al.
-2020), the true potential of these models lies in their application to high-quality private data.
-Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (Yang et al.
-2024).
-Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (Ye et al.
-2024).
+While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (kaplan2020scaling), the true potential of these models lies in their application to high-quality private data.
+Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (yang2024federated).
+Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (ye2024openfedllm).
 Therefore, effectively incorporating private data will further advance the performance of LLMs.
@@
 The distributed private data silos and the significant computational demands of LLMs pose substantial challenges for their deployment in federated environments.
-Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprise servers (Huang et al.
-2024), or medical institutions (Yuan et al.
-2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al.
-2024).
-Federated Learning (FL) (Stremmel and Singh 2021) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation.
-They often leverage Adapters (Houlsby et al.
-2019) or LoRA (Hu et al.
-2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to fine-tune LLMs.
+Private data is often isolated in local silos, such as mobile devices (ghimire2022recent), enterprise servers (huang2024federated), or medical institutions (yuan2024decentralized), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (chen2024federated).
+Federated Learning (FL) (stremmel2021pretraining) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation.
+They often leverage Adapters (houlsby2019parameter) or LoRA (hu2022lora) as transmitted parameters, or employ collaborative training between large and small models (hinton2015distilling) to fine-tune LLMs.
 However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible.
 Therefore, how to effectively deploy LLMs in FL frameworks while ensuring privacy and scalability in resource-constrained scenarios remains an open research problem.
-To address this, transformer-based federated splitting models (Zheng et al.
-2024; Su et al.
-2024) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios.
-Titanic (Su et al.
-2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training.
-FL-GLM (Zheng et al.
-2024) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client.
+To address this, transformer-based federated splitting models (zheng2024safely; su2024titanic) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios.
+Titanic (su2024titanic) handles limited server resources by partitioning LLM across clients and enabling collaborative training.
+FL-GLM (zheng2024safely) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client.
 As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on the client, while the middle blocks on the server.
@@
 After the client calculates loss, gradients backpropagate (BP) from client-to-server-to-client with the same encryption.
-Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al.
-2023).
+Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (asnani2023reverse).
 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens.
@@
 Federated Learning for LLMs
-Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (McMahan et al.
-2017; Stremmel and Singh 2021; Ji et al.
-2019).
-Recent federated LLM frameworks (Fan et al.
-2023; Kuang et al.
-2024; Ye et al.
-2024) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead.
+Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (mcmahan2017communication; stremmel2021pretraining; ji2019learning).
+Recent federated LLM frameworks (fan2023fate; kuang2024federatedscope; ye2024openfedllm) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead.
 However, transmission demands often still exceed system limits.
-To mitigate this, gradient compression (Wu et al.
-2024b; Shu et al.
-2024) and quantization (JianHao et al.
-2024) techniques have been applied.
-Structural pruning based on Fisher Information (Liu et al.
-2024) further narrows updates to key layers, reducing model size without performance loss.
-Knowledge distillation (Wu et al.
-2024a; Fan et al.
-2024, 2025) offers another solution, using smaller proxy models to transfer knowledge with minimal communication.
-Client heterogeneity remains a major challenge (Zhang et al.
-2024), prompting methods like LoRA-based stacking (Bai et al.
-2024; Wang et al.
-2024) to align models from diverse clients and improve fairness.
+To mitigate this, gradient compression (wu2024cg; shu2024ferret) and quantization (jianhao2024promoting) techniques have been applied.
+Structural pruning based on Fisher Information (liu2024fisher) further narrows updates to key layers, reducing model size without performance loss.
+Knowledge distillation (wu2024fedbiot; fan2024fedcollm; fan2025fedmkt) offers another solution, using smaller proxy models to transfer knowledge with minimal communication.
+Client heterogeneity remains a major challenge (zhang2024towards), prompting methods like LoRA-based stacking (bai2024federated; wang2024flora) to align models from diverse clients and improve fairness.
 Despite advances in efficient fine-tuning and parameter reduction for federated LLMs, resource-constrained clients still face challenges with independent fine-tuning, particularly under compression overhead.
-Split learning (Thapa et al.
-2022) offers a promising alternative in such settings.
+Split learning (thapa2022splitfed) offers a promising alternative in such settings.
 Split Learning for LLMs
-Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al.
-2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients.
+Split learning is an emerging decentralized training approach designed for resource-constrained settings (abedi2024fedsl; rahman2020internet; matsubara2021neural), especially mobile or GPU-limited clients.
 It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning.
-To address the training of sequential data in LLMs, FedBERT (Tian et al.
-2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices.
+To address the training of sequential data in LLMs, FedBERT (tian2022fedbert) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices.
 Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points, and susceptibility to privacy threats like embedding gradient attacks.
-Recently, Su et al.
-(2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity.
+Recently, su2024titanic extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity.
 In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation.
@@
 Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities.
-Zheng et al.
-(2024) propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources.
+zheng2024safely propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources.
 And then, they employ key encryption during client-server communication to prevent reverse engineering attacks from peer clients.
@@
 We design an extreme attack scenario to evaluate the robustness of FedSEA-LLaMA.
-Inspired by Pasquini, Ateniese, and Bernaschi (2021), we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server.
+Inspired by pasquini2021unleashing, we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server.
 The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally.
@@
 where $\text{RMSNorm}(*)$ is the layernorm function, $\mathcal{A}$ is the attention layer, $\mathcal{M}$ is the MLP layer.
-During backpropagation, according to (Clark 2017), the gradient of $W_{n}$ is:
+During backpropagation, according to (clark2017computing), the gradient of $W_{n}$ is:
 $\frac{\partial L}{\partial W_{n}}$ $=\frac{\partial L}{\partial h_{5}}\cdot\frac{\partial h_{5}}{\partial h_{4}}\cdot\frac{\partial h_{4}}{\partial W_{n}},$ (8)
@@
 Acc.
-F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al.
-2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al.
-2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al.
-2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al.
-2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (Zheng et al.
-2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (Touvron et al.
-2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
+F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (raffel2020exploring) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (lewis2020bart) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (glm) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (zeng2022glm) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (zheng2024safely) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (touvron2023LLaMA) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
 Table 1: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FedSEA-LLaMA and baselines.
@@
 To evaluate model generalization across diverse NLP tasks, our experiments incorporate three distinct tasks:
-  • The SuperGLUE benchmark (Wang et al. 2020) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
-  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains.
+  • The SuperGLUE benchmark (wang2020superglue) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
+  • The CoQA dataset (reddy2019coqa) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains.
 Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
-  • The XSum dataset (Narayan, Cohen, and Lapata 2018) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries.
+  • The XSum dataset (narayan2018don) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries.
 It tests the model’s ability to abstract and condense salient information across diverse topics.
@@
 For NLU tasks such as COPA, WSC, etc, we report Accuracy, which measures the proportion of instances where the model’s predicted label exactly matches the ground-truth label.
-For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (Yatskar 2019; You et al.
-2022).
+For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (yatskar2019qualitative; you2022end).
 EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
-For summarization tasks, we use standard ROUGE metrics (Lin 2004; Liu et al.
-2021; Fang et al.
-2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
+For summarization tasks, we use standard ROUGE metrics (lin2004rouge; liu2021topic; fang2022spoken), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
 Baselines
-We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (Touvron et al.
-2023), an open-source pre-trained model developed by Meta and FL-GLM.
-We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (Raffel et al.
-2020), $\text{BART}_{\text{Large}}$ (Lewis et al.
-2020), $\text{GLM}_{\text{RoBERTa}}$ (Du et al.
-2022) and ChatGLM-6B (GLM et al.
-2024).
+We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (touvron2023LLaMA), an open-source pre-trained model developed by Meta and FL-GLM.
+We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (raffel2020exploring), $\text{BART}_{\text{Large}}$ (lewis2020bart), $\text{GLM}_{\text{RoBERTa}}$ (glm) and ChatGLM-6B (glm2024chatglm).
 Parameter Settings
 The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40.
-The implementation is based on the Flower (Beutel et al.
-2020) framework, which orchestrates the transfer of intermediate data between clients and the server.
+The implementation is based on the Flower (beutel2020flower) framework, which orchestrates the transfer of intermediate data between clients and the server.
 Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset.
@@
 62406013,No.62388101, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-24-034), and the Fundamental Research Funds for the Central Universities.
-Appendix A Noise Added to Hidden States
-(a) Loss Curves for Noise Scales
-($Blocks_{A}$=1, $Blocks_{C}$=1)
-(b) Loss Curves for Noise Scales
-($Blocks_{A}$=2, $Blocks_{C}$=1)
-(c) Loss Curves for Noise Scales
-($Blocks_{A}$=3, $Blocks_{C}$=1)
-Figure 6: FedSEA-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
-The magnitude of noise injection significantly influences the stability and convergence speed of model training.
-Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance.
-To quantify this effect, we first fine-tune the FedSEA-LLaMA model on the MultiRC (Khashabi et al.
-2018) with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within the hidden state $h_{A}$.
-According to Figure 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1).
-The average magnitude of the hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$.
-Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance.
-Training is conducted for 5,000 steps with FP16 precision.
-We use LoRA with a rank of 8 and a scaling factor $\alpha=16$.
-The learning rate is 1e-5, and the batch size is 2.
-Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2.
-To assess the sensitivity to the noise scale when $Blocks_{A}$=3, we evaluate performance across a range of noise scales (from 0.1 to 0.4).
-As shown in Table 5, model performance decreases as the noise scale increases.
-This suggests that while moderate noise injection is tolerable, excessive noise harms representation quality.
-Noise Scale F1a EM
-0.1 75.68 34.10
-0.15 71.21 26.82
-0.2 59.08 16.47
-0.3 55.79 7.45
-0.4 57.68 6.09
-Table 5: Performance under different noise scales when $Blocks_{A}$=3
-A horizontal comparison reveals that, under the same noise level, increasing the number of blocks in $Blocks_{A}$ consistently accelerates convergence.
-Table 2 indicates that when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.
-This can be explained by the average magnitude shown in Figure 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR).
-This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
-Figure 7: Avg magnitude of hidden state $h_{A}$ ($Blocks_{C}$ =1) on MultiRC dataset.
-Appendix B Prompt
-Table 6 illustrates examples of cloze-style questions and corresponding answers derived from each task.
-(1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages.
-(2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise.
-(3)WSC focuses on pronoun disambiguation requiring deep contextual understanding.
-(4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another.
-(5)BoolQ involves yes/no questions answered using evidence from context passages.
-(6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences.
-(7)CB deals with entailment in complex sentences with nuanced reasoning.
-(8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information.
-Table 7 illustrates examples of fine-tuning templates of CoQA and XSum.
-Dataset Task Cloze Question Answers ReCoRD Question answering [passage p] [cloze question q] Answer candidates COPA Causal reasoning “[choice c1]” or“[choice c2]”?
-[premise p], so [M].
-c1/c2 WSC Coreference resolution [sentence s] The pronoun ‘*p*’ refers to [M].
-Noun n RTE Textual entailment “[hypothesis h]”?
-[M] “[premise p]” “yes”/“no” BoolQ Question answering [passage p].
-Question: q?
-Answer:[M].
-“yes” / “no” WiC Word sense disambiguation “[sentence s1]”/“[sentence s2]”Similar sense of [word w]?
-[M].
-“yes”/“no” CB Textual entailment “[hypothesis h]”?
-[M], “[premise p]” “yes”/“no”/“maybe” MultiRC Question answering [passage p].
-Question: q?
-Is it [answer a]?
-[M].
-“yes”/“no”
-Table 6: Cloze questions and answers for the 8 SuperGLUE tasks
-Dataset Task Fine-tuning template
-CoQA conversational QA
-<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
-You need to answer my question base on the provided story.<</SYS>><</SYS>>
-The story is {story}.My question is {question}.[/INST]{answer}</s></s>
-XSum summarization generation
-<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
-You need to summarize an abstract base on the provided passage.<</SYS>><</SYS>>
-The passage is {passage}.[/INST]{summarization} </s></s>
-Table 7: Fine-tuning templates for CoQA and XSum
-Appendix C Parallel Training
-[…]
-Table 8: Comparison of training time between different training strategies on ReCoRD dataset.
-(a) Client-batch parallel training.
-(b) Server-hierachical parallel training.
-Figure 8: Multi-client Parallel Training of FedSEA-LLaMA.
-In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation.
-Then, the server splits hidden state $h_{B}$ and sends them back to different clients.
-In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$.
-After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
-In multi-client settings, the FedSEA-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism, and server-hierarchical parallelism, each designed for different system configurations and resource availabilities.
-Sequential Training
-The client begins by passing hidden states through its input blocks $Blocks_{A}$ and forwards the result to the server.
-The server processes the input through $Blocks_{B}$ and returns the output states.
-The client then computes the loss with $Blocks_{C}$, sends the gradients to the server for model updates, and receives the updated gradients to adjust its input blocks.
-This process continues until all clients have completed their respective training phases for the current round.
-Although this method is inherently slower due to its serialized nature, it significantly reduces the computational burden on the server, making it a suitable choice for scenarios where server-side resources are limited.
-Client-batch Parallelism
-The client-batch parallel approach enhances training throughput by expanding the effective batch size at the server, as shown in Figure  8(a).
-During each training round, the server collects intermediate hidden states from all participating clients.
-If each of the $M$ clients processes a local batch of size $b$, the hidden states sent to the server will each have the shape $(b,\textit{seq\_length},\textit{hidden\_size})$.
-The server concatenates these hidden states along the batch dimension, forming a combined tensor of shape $(b\times M,\textit{seq\_length},\textit{hidden\_size})$.
-This aggregated batch is then used to perform a joint forward and backward pass on the server model, allowing it to compute gradients more efficiently and update shared parameters.
-Server-hierarchical Parallelism
-In a server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Figure  8(b).
-These instances run in parallel, enabling simultaneous local training for all clients.
-Typically, after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers.
-This is usually done via weighted averaging of both server-side and client-side model parameters.
-This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
-For communication cost analysis, we randomly select 1,000 samples from the ReCoRD dataset.
-All training uses a batch size of 1 with FP16 precision for efficiency.
-LoRA is applied with a rank of 8 and a scaling factor of 16, targeting attention modules during fine-tuning.
-The model is partitioned with $Blocks_{A}$=1 and $Blocks_{C}$=1.
-Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time in Table 8.
-Results show that sequential training is the slowest, client-batch training is faster, and server hierarchical training is the fastest.
-Notably, when the server can handle larger batches, increasing the number of clients in client-batch training effectively reduces training time.
-Appendix D Efficiency Comparison between FedSEA-LLaMA and FL-GLM
-For the measurements of inference speed, inference memory usage, and per-token communication volume, we used the following prompt template:
-{B_INST}{B_SYS}System: You are a helpful assistant. Always answer as helpfully as possible.{E_SYS}User: what can you do?{E_INST}
-The total prompt length was 44 tokens.
-All inference experiments were conducted in FP16 precision with a maximum output length of 128 tokens.
-During inference, the client held one layer of model A and one layer of model C.
-For the training memory usage experiments, we used the MultiRC dataset.
-Both model A and model C contained one layer.
-Unless otherwise specified, the default configuration was as follows: learning rate $1\times 10^{-5}$, LoRA rank =32=32, LoRA scaling factor =32=32, random seed =64=64, FP16 precision, and batch size =2=2.
-It is important to note that our training implementation actively clears unused variables in real time to minimize GPU memory usage.
-Therefore, the reported memory values correspond to the maximum observed usage within a one-minute monitoring window.
-All experiments were repeated three times with different random seeds (42, 64, and 128) to ensure the robustness and reproducibility of the results.
-The comparison results are shown in the Table 9.
-Framework
-Training Memory Usage
-Inference Memory Usage
-per-token communication
-volume (Bytes)
-Inference tokens
-per second
-FL-GLM 9810±0.00 6886±0.00 context length N × 10,335 0.75±0.03
-FedSEA-LLaMA 5078±0.00 3674.00±0.00 8494 0.34±0.01
-Table 9: Efficiency Comparison between FedSEA-LLaMA and FL-GLM
-Appendix E Parameter Setting
-FedSEA-LLaMA and Baselines
-Different hyperparameters are explored to obtain the best performance of FedSEA-LLaMA and Centralized LLaMA.
-Specifically, we tune the LoRA rank (r) and scaling factor (alpha) as pairs, trying combinations including (4, 32), (8, 32), (16, 32), (32, 32), and (8, 16).
-For the learning rate (lr), values of 1e-5, 2e-5, and 3e-5 are tested.
-The final parameter settings were selected based on the model’s validation performance.
-The final parameters are shown in Table 10.
-CoQA XSum MultiRC Record
-FedSEA-LLaMA Lora r & alpha $(4,32)$ $(16,32)$ $(32,32)$ $(32,32)$
-lr 1e-4 3e-5 1e-5 2e-5
-Centralized-LLaMA Lora r & alpha $(8,16)$ $(32,32)$ $(32,32)$ $(8,32)$
-lr 5e-5 3e-5 1e-5 3e-5
-Table 10: Hyperparameters for FedSEA-LLaMA and Baselines
-Different Noise Scales and Layers of $Block_{A}$ Experiment
-We use a LoRA rank of 8 with a scaling factor of 16, a learning rate of 1e-5, and FP16 precision to investigate the effects of different noise scales and the number of blocks in $Blocks_{A}$ on downstream performance.
-Ablation Study on Attention-Mask
-The context length is set to 500, and the precision is set to FP16.
-Ablation Study on KV Cache
-The precision is set to FP16. Different context lengths are set to investigate the inference speed of FedSEA-LLaMA and FL-GLM.
-Different Model Partition
-All experimental results were obtained after 8,000 training steps with LoRA rank set to 8, scaling factor set to 16, a batch size of 1, FP16 precision, and a learning rate of 1e-5.